### v1

# Generic graded contractions of Lie algebras

Generic graded contractions of Lie algebras
Mikhail V. Kochetov and Serhii D. Koval†‡
Department of Mathematics and Statistics, Memorial University of Newfoundland,
  St. John’s (NL) A1C 5S7, Canada
Institute of Mathematics of NAS of Ukraine, 3 Tereshchenkivska Str., 01024 Kyiv, Ukraine
E-mails: mikhail@mun.ca, skoval@mun.ca
We study generic graded contractions of Lie algebras from the perspectives of group cohomology, affine algebraic geometry and monoidal categories. We show that generic graded contractions with a fixed support are classified by a certain abelian group, which we explicitly describe. Analyzing the variety of generic graded contractions as an affine algebraic variety allows us to describe which generic graded contractions define graded degenerations of a given graded Lie algebra. Using the interpretation of generic $G$-graded contractions as lax monoidal structures on the identity endofunctor of the monoidal category of $G$-graded vector spaces, we establish a functorial version of the Weimar–Woods conjecture on equivalence of generic graded contractions.
1 Introduction
Deformation theory of different algebraic and geometric objects has been a subject of significant interest in mathematics ever since the study of Riemann surfaces and their moduli spaces by Riemann in XIX century, see a review of the history of the subject in [15]. There are various types of deformations appearing in algebra: one-parameter deformations, degenerations and contractions. From the perspective of physics, the most useful deformations are those that connect symmetry groups (or their Lie algebras) when one physical theory can be obtained as a limit of another. For example, Poincaré group, which is the symmetry group of relativistic mechanics, ‘transforms’ into Galilean group, the symmetry group of classical mechanics, when the speed of light is assumed to approach infinity. This example motivated Segal [35] to initiate the study of contractions of Lie groups and Lie algebras. Later, Inönü and Wigner [22, 23] considered the simplest types of one-parameter continuous contractions, where some of the basis elements of the Lie algebra are scaled by a parameter that tends to zero. This type of contractions is now called the Inönü–Wigner contractions. For example, the transition from Heisenberg algebra to abelian algebra of the same dimension, which represents the transition from quantum mechanics to classical mechanics when Planck’s constant goes to zero, is an Inönü–Wigner contraction. Another famous case of an Inönü–Wigner contraction is the transition from de Sitter invariant relativity to special relativity, corresponding to a contraction from de Sitter algebra to Poincaré algebra.
Despite their simplicity, Inönü–Wigner contractions found a wide range of applications in both mathematics and physics. Saletan [34] and then Doebner and Melsheimer [10] generalized this concept in several directions, leading to what are now known as generalized Inönü–Wigner contractions. These generalized contractions provide a more convenient framework for applications to the study of representations, invariants and special functions, see, for example, [37]. For a more comprehensive literature review on this topic, as well as the state of the art on contractions between low-dimensional real and complex Lie algebras, see, e.g., [31]. Contractions of some infinite-dimensional Lie algebras (Witt, Virasoro and affine Kac–Moody algebras) were studied in [16, 17].
Contractions of Lie algebras are closely related to degenerations. In fact, over the field of complex numbers these two concepts coincide due to Borel’s closed orbit lemma. Milnor, in his seminal work on invariant metrics on Lie groups [26], implicitely used degenerations of Lie algebras to demonstrate the existence of invariant Rimannian metrics on Lie groups.
Essentially, degenerations of an $n$-dimensional Lie algebra $\mathfrak{L}$ are limit points of its ${\rm GL}_{n}(\mathbb{F})$-orbit $\mathcal{O}(\mathfrak{L})$ in the variety of $n$-dimensional Lie algebras endowed with the Zariski topology. When the orbit $\mathcal{O}(\mathfrak{L})$ is a Zariski open set, the Lie algebra $\mathfrak{L}$ is called rigid. Intuitively, a Lie algebra is rigid if all Lie algebras close to it (in the sense of Zariski topology) are isomorphic. Rigidity of Lie algebras has been extensively studied and characterized in different terms, see, e.g., [5, 18] and references therein. The classification of rigid Lie algebras remains an important open problem.
While the general linear group ${\rm GL}_{n}(\mathbb{F})$ is the standard choice to study degenerations, in the context of Lie algebras graded by an abelian group $G$, it is natural to replace it with an algebraic torus, see Subsection 2.2. Montigny and Patera in [30] defined a new type of Lie algebra contractions, which are in general multi-parametric and not necessarily continuous. According to the authors, this provides “a much larger variety of contraction ‘limits’ to be studied”. This type of contractions is known as graded contractions, and its defining feature is the preservation of a chosen $G$-grading during the contraction process. In Section 4, we investigate which among these contractions are graded degenerations (Corollary 30) and, in particular, related to Inönü–Wigner contractions.
Over the past three decades, graded contractions have been extensively studied from various perspectives, including computations for specific graded Lie algebras [3, 6, 8, 9, 20, 21, 25, 29] and theoretical aspects [28, 39, 40]. The concept of graded contraction is flexible with respect to changes in the underlying category of algebras and thus can be easily transfered to other structures, such as superalgebras [30] and Jordan algebras [24]. For a review of the state of the art in graded contractions, we refer to the recent paper by Draper, Sánchez-Ortega and Meyer [9].
For generic graded contractions, which are studied in [28, 37, 39] and are also the subject of the present paper, the defining conditions are chosen in such a way that the contraction can be applied uniformly to all $G$-graded Lie algebras. This allows us to study such contractions from various perspectives: group cohomology, affine algebraic geometry and monoidal categories. As a result, we not only generalize and reinterpret existing results on classification of such contractions from [28, 39], but also place them into a wider theoretical framework.
The structure of the paper is as follows. In Section 2, theoretical background on gradings, contractions, degenerations, and graded contractions as well as their equivalences are reviewed.
In Section 3, we discuss the classification of generic graded contractions. To any subset $S$ of the set of unordered pairs in $G$, we associate free abelian groups of “surviving defining relations” and “surviving higher-order identities” and use them to show that contractions with support $S$ are classified by a certain abelian group $H^{2}_{S}(\mathbb{F}^{\times})$, which can be computed using the exact sequences presented in Theorem 19. In particular, for an algebraically closed or a real closed field $\mathbb{F}$, we reinterpret the results of Weimar–Woods [39].
Section 4 is devoted to the study of the affine variety of generic graded contractions. We study the action of an algebraic torus on this variety and the orbit closures of contractions. In Theorem 27, we describe the vanishing ideal of such an orbit closure. We show that the contractions in the closure of the unique open orbit are precisely the “continuous contractions” in the sense of [28] and relate them to Inönü–Wigner contractions using Theorem 19 and [38, Theorem VI.2].
In Section 5, we discuss the Weimar–Woods conjecture that two graded contractions are strongly equivalent if and only if they are equivalent via normalization. We show that a generic $G$-graded contraction endows the identity endofunctor of the monoidal category of $G$-graded vector spaces with a lax monoidal structure. This allows us to prove a functorial version of the Weimar–Woods conjecture: two graded contractions define naturally isomorphic lax monoidal endofunctors if and only if they are equivalent via normalization (Proposition 36). In Section 6, we summarize the results of the paper, discuss open problems and possible directions for future research.
2 Preliminaries
We review some basic facts about contractions, gradings and graded contractions of Lie algebras following [9, 13, 30, 31].
2.1 Contractions and degenerations of Lie algebras
Contractions. Let $\mathfrak{L}:=(V,[\cdot,\cdot])$ be a real or complex Lie algebra with an underlying vector space $V$ and a Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]. Consider a continuous function $U\colon(0,1]\to{\rm GL}(V)$. A parameterized family of new Lie brackets on $V$ is defined in terms of the Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] in the following way:
$\forall t\in(0,1],\quad\forall x,y\in V\colon\qquad[x,y]_{t}:=U_{t}^{-1}[U_{t}x,U_{t}y].$ (1)
It is clear that for any $t\in(0,1]$ the Lie algebra $\mathfrak{L}_{t}:=(V,[\cdot,\cdot]_{t})$ is isomorphic to $\mathfrak{L}$.
Definition 1.
If the limit $[x,y]_{0}:=\lim_{t\to 0^{+}}[x,y]_{t}$ exists for any $x,y\in V$, then $[\cdot,\cdot]_{0}$ is a well-defined Lie bracket on $V$ and the Lie algebra $\mathfrak{L}_{0}:=(V,[\cdot,\cdot]_{0})$ is called a one-parametric contionuous contraction (or simply contraction) of the Lie algebra $\mathfrak{L}=(V,[\cdot,\cdot])$.
Among the functions $U\colon(0,1]\to{\rm GL}(V)$ those realized by specific diagonal matrices are particularly important for the study of contractions and their applications in mathematical physics. Consider the following family of such functions:
$U_{t}=\bigoplus_{s=0}^{m}t^{n_{s}}{\rm Id}_{V^{(s)}},\quad\mbox{where}\quad V=\bigoplus_{s=0}^{m}V^{(s)}\quad\mbox{and}\quad n_{0},\dots,n_{m}\in\mathbb{Z},\quad m\geqslant 1.$ (2)
Choosing a basis in each subspace $V^{(s)}$, we represent $U_{t}$ by a diagonal matrix with entries $t^{n_{s}}$, each repeated $\dim V^{(s)}$ times. Then the limit $[\cdot,\cdot]_{0}$ exists if and only if the following constraints are satisfied:
$n_{s_{i}}+n_{s_{j}}\geqslant n_{s_{k}},\quad\mbox{whenever}\quad c_{ij}^{k}\neq 0,\quad i,j,k=1,\dots,\dim V,$ (3)
where $c_{ij}^{k}$ stands for the components of the structure tensor of the Lie algebra $\mathfrak{L}$ in the chosen basis, with the $i$-th basis element belonging to $V^{(s_{i})}$. When these conditions hold, the contracted algebra $\mathfrak{L}_{0}$ exists and is referred as a generalized Inönü–Wigner contraction of $\mathfrak{L}$. In the particular case $m=1$, $n_{0}=0$ and $n_{1}=1$, the contracted algebra $\mathfrak{L}_{0}$ is called a simple Inönü–Wigner contraction.
Remark 2.
As noted in [19], we can associate to the decomposition $V=\bigoplus_{s=0}^{m}V^{(s)}$ in (2) the filtration
$\mathfrak{L}:=\bigcup_{n\in\mathbb{Z}}\mathfrak{L}(n)\quad\mbox{with}\quad\mathfrak{L}(n)=\bigoplus_{i\in\{0,\dots,m\},\ n_{i}\leqslant n}V^{(i)}.$
The constraints (3) are necessary and sufficient to ensure that $[\mathfrak{L}(n),\mathfrak{L}(m)]\subset\mathfrak{L}(n+m)$. The associated graded algebra $\operatorname{gr}\mathfrak{L}$ of $\mathfrak{L}$ with respect to the above filtration is isomorphic to the contracted algebra $\mathfrak{L}_{0}$ under the contraction matrix $U(\varepsilon)$ from (2). This follows directly from the description of the structure tensor $\tilde{c}_{ij}^{k}$ of $\mathfrak{L}_{0}$, see, e.g., [31, Section 3]:
$\tilde{c}_{ij}^{k}=c_{ij}^{k}\quad\mbox{if}\quad n_{s_{i}}+n_{s_{j}}=n_{s_{k}}\quad\mbox{and}\quad\tilde{c}_{ij}^{k}=0\quad\mbox{otherwise}.$
Degenerations. Consider an $n$-dimensional vector space $V$ over a field $\mathbb{F}$. Denote by $\mathcal{L}_{n}(\mathbb{F})$ the variety of Lie algebra laws on $V$, i.e., the set of all alternating bilinear maps from $V\times V$ to $V$ satisfying the Jacobi identity. The set $\mathcal{L}_{n}(\mathbb{F})$ is an algebraic set in the affine space $\wedge^{2}V^{*}\otimes V$, which is invariant under the natural action of ${\rm GL}(V)$ on $\wedge^{2}V^{*}\otimes V$, given by
$(g\cdot\mu)(x,y)=g(\mu(g^{-1}x,g^{-1}y))\quad\mbox{for all}\ \ g\in{\rm GL}(V),\ \mu\in\mathcal{L}_{n}(\mathbb{F})\ \mbox{and}\ x,y\in V.$
(Note that this is a left action, whereas the one in (1), following the tradition in physics literature, is on the right.) Denote by $\mathcal{O}(\mu)$ the orbit of $\mu\in\mathcal{L}_{n}(\mathbb{F})$ under the action of ${\rm GL}(V)$ and by $\overline{\mathcal{O}(\mu)}$ its closure with respect to the Zariski topology in $\mathcal{L}_{n}(\mathbb{F})$.
Definition 3.
The Lie algebra $\mathfrak{L}_{0}:=(V,\mu_{0})$ is called a degeneration of the Lie algebra $\mathfrak{L}:=(V,\mu)$ if $\mu_{0}\in\overline{\mathcal{O}(\mu)}$.
When $\mathbb{F}$ is algebraically closed, the orbit $\mathcal{O}(\mu)$ is open in its Zariski closure $\overline{\mathcal{O}(\mu)}$ by the closed orbit lemma in [4]. Hence, the boundary $\overline{\mathcal{O}(\mu)}\setminus\mathcal{O}(\mu)$ is a union of orbits of strictly lower dimension. Moreover, all points of $\mathcal{O}(\mu)$ are regular in $\overline{\mathcal{O}(\mu)}$. Finally, the orbit $\mathcal{O}(\mu)$ is a constructible set, so in the case $\mathbb{F}=\mathbb{C}$, the closure of $\mathcal{O}(\mu)$ in Zariski topology coincides with its closure in the metric topology, see [27, Corollary 1, p. 60]. Therefore, over $\mathbb{C}$, the concepts of contraction and degeneration coincide.
It is worth pointing out that, over $\mathbb{R}$, any contraction is a degeneration. The converse probably fails because the metric closure of an orbit may be strictly contained in the Zariski closure, see examples in [11].
2.2 Gradings and graded contractions
Let $G$ be an abelian group and $\mathbb{F}$ be a field. A $G$-grading on an arbitrary $\mathbb{F}$-algebra $A$ (not necessarily unital or associative) is a vector space decomposition $\Gamma\colon A=\bigoplus_{g\in G}A_{g}$ such that $A_{g}A_{h}\subseteq A_{gh}$ for all $g,h\in G$. The elements of $A_{g}$ are said to be homogeneous of degree $g$ and the subspaces $A_{g}$ are called homogeneous components. The support of the grading $\Gamma$ is the set ${\rm Supp}(\Gamma):=\{g\in G\colon A_{g}\neq 0\}$. If $\Gamma$ is fixed, $(A,\Gamma)$ is said to be a $G$-graded algebra. We may write $A$ for a graded algebra if the grading is clear from the context. For more detailed exposition see, e.g., [13].
Graded contractions. Denote the multiplication of an algebra $A$ by $\mu$. Given an arbitrary map $\varepsilon\colon G\times G\to\mathbb{F}$, we can define a new multiplication on the underlying vector space of $A$:
$\mu^{\varepsilon}(x,y)=\varepsilon(g,h)\mu(x,y)\quad\mbox{for all}\ \ g,h\in G,\ x\in A_{g},\ y\in A_{h},$
which defines a new algebra, $A^{\varepsilon}$. However, if $A$ belongs to a specific variety (e.g., associative, commutative, Lie or Jordan algebras), $A^{\varepsilon}$ is not guaranteed to belong to the same variety unless we impose restrictions on $\varepsilon$. For example, if $\varepsilon$ is a two-cocycle on $G$ with values in $\mathbb{F}^{\times}$, then $A^{\varepsilon}$ is known as a cocycle twist of $A$. In this case, if $A$ is an associative algebra, then so is $A^{\varepsilon}$; also $(A^{\varepsilon})^{\varepsilon^{-1}}=A$. In the context of Lie algebras, the definition of a graded contraction takes the following form:
Definition 4.
Let $\Gamma\colon\mathfrak{L}=\bigoplus_{g\in G}\mathfrak{L}_{g}$ be a $G$-grading on a Lie algebra $\mathfrak{L}$. A graded contraction of $(\mathfrak{L},\Gamma)$ is a map $\varepsilon\colon G\times G\to\mathbb{F}$ such that $\mathfrak{L}^{\varepsilon}$ is a Lie algebra. Then $\mathfrak{L}^{\varepsilon}$ is called the contracted algebra or, by abuse of language, the graded contraction of $(\mathfrak{L},\Gamma)$ with respect to $\varepsilon$.
The algebra $\mathfrak{L}^{\varepsilon}$ is a Lie algebra if and only if the following conditions are satisfied:
  1. $(i)$
    $\big{(}\varepsilon(g,h)-\varepsilon(h,g)\big{)}[x,y]=0$,
  2. $(ii)$
    $\big{(}\varepsilon(g,hk)\varepsilon(h,k)-\varepsilon(k,gh)\varepsilon(g,h)\big{)}[x,[y,z]]+\big{(}\varepsilon(h,kg)\varepsilon(k,g)-\varepsilon(k,gh)\varepsilon(g,h)\big{)}[y,[z,x]]=0$
for all $x\in\mathfrak{L}_{g}$, $y\in\mathfrak{L}_{h}$ and $z\in\mathfrak{L}_{k}$. Conditions $(i)$ and $(ii)$ arise from the antisymmetry and the Jacobi identity, respectively, applied to $\mathfrak{L}^{\varepsilon}$.
Fixing a Lie algebra $\mathfrak{L}$ and a $G$-grading $\Gamma$ on it, the investigation of graded contractions of $(\mathfrak{L},\Gamma)$ can be quite laborious [20, 21]. Trivially, when $\mathfrak{L}_{gh}=0$ and $\mathfrak{L}_{ghk}=0$, conditions $(i)$ and $(ii)$ are satisfied automatically, imposing no constraints on the values $\varepsilon(g,h)$.
To develop a general theory of graded contractions, it is natural to disregard the specifics of the grading on $\mathfrak{L}$ and consider the so-called “generic case”, i.e., require that $(i)$ and $(ii)$ hold for all $G$-graded Lie algebras.
Proposition 5.
The necessary and sufficient conditions for the map $\varepsilon\colon G\times G\to\mathbb{F}$ to define a $G$-graded contraction for all $G$-graded Lie algebras are
  1. $(i^{\prime})$
    $\varepsilon(g,h)=\varepsilon(h,g)$,
  2. $(ii^{\prime})$
    $\varepsilon(g,hk)\varepsilon(h,k)=\varepsilon(k,gh)\varepsilon(g,h)$
for all $g,h,k\in G$.
Proof.
Since $(ii^{\prime})$ implies $\varepsilon(k,gh)\varepsilon(g,h)=\varepsilon(h,kg)\varepsilon(k,g)$ by renaming variables, it is clear that $(i^{\prime})$ and $(ii^{\prime})$ are sufficient for $\varepsilon$ to define a graded contraction for any $(\mathfrak{L},\Gamma)$.
To prove the necessity, it suffices to find a $G$-graded Lie algebra $\mathfrak{L}$ satisfying (a) $[\mathfrak{L}_{g},\mathfrak{L}_{h}]\neq 0$ for all distinct $g,h\in G$, (b) for all $g,h,k\in G$ there exist $x\in\mathfrak{L}_{g}$, $y\in\mathfrak{L}_{h}$, $z\in\mathfrak{L}_{k}$ such that $[x,[y,z]]$ and $[y,[z,x]]$ are linearly independent. For instance, consider the free Lie algebra $\mathfrak{F}$ on the generators $x_{g}$, $y_{g}$ and $z_{g}$, where $g$ runs through $G$. This algebra has a natural grading by setting the $G$-degree of generators $x_{g}$, $y_{g}$ and $z_{g}$ to be $g$. Since $[x_{g},x_{h}]\neq 0$ for all $g\neq h$, condition $(i)$ implies $(i^{\prime})$ for the graded Lie algebra $\mathfrak{F}$. Since the commutators $[x_{g}[y_{h},z_{k}]]$ and $[y_{h}[z_{k},x_{g}]]$ are linearly independent in $\mathfrak{F}$, condition $(ii)$ implies $(ii^{\prime})$. ∎
Remark 6.
If $G$ is finite, then we can restrict to finite-dimensional Lie algebras in the above proposition. Indeed, we can replace the Lie algebra $\mathfrak{F}$ with the finite-dimensional nilpotent Lie algebra $\mathfrak{F}/\mathfrak{F}^{4}$, where $\mathfrak{F}^{4}$ is the fourth term of the lower central series of $\mathfrak{F}$.
When $\varepsilon$ does not take value 00, conditions $(i^{\prime})$ and $(ii^{\prime})$ say that $\varepsilon$ is a symmetric two-cocycle with values in $\mathbb{F}^{\times}$ with trivial $G$-action: $\varepsilon\in Z^{2}_{sym}(G,\mathbb{F}^{\times})$.
In what follows, we restrict our attention to this generic point of view, as was done, for example, in [28, 39]. All definitions and examples in this paper are presented within this framework. In particular, when using the notion “graded contraction” we omit referring to a $G$-graded Lie algebra $(\mathfrak{L},\Gamma)$. For a “grading-dependent” treatment of contractions, see, for example, [8, 9, 30].
The obvious examples of the graded contractions are the constant maps $\varepsilon=1$ and $\varepsilon=0$. Their corresponding contracted algebras are respectively the original algebra $\mathfrak{L}$ and the abelian Lie algebra. Both these maps are examples of trivial graded contractions, that is, the graded contractions $\varepsilon\colon G\times G\to\mathbb{F}$ such that the for all $G$-graded Lie algebras $\mathfrak{L}$ the contracted algebra $\mathfrak{L}^{\varepsilon}$ is isomorphic to $\mathfrak{L}$ or to the abelian Lie algebra on the vector space $\mathfrak{L}$, see [39, Definition 2.6]
A more interesting example of a trivial graded contraction is that arising from the normalization of homogeneous components of the grading $\Gamma$. Let $\alpha\colon G\to\mathbb{F}^{\times}$ be an arbitrary map, then the map $d\alpha\colon G\times G\to\mathbb{F}^{\times}$,
$d\alpha(g,h)=\frac{\alpha(g)\alpha(h)}{\alpha(gh)},\quad\forall g,h\in G,$
is a graded contraction of $\mathfrak{L}$. It is clear that $\mathfrak{L}^{d\alpha}$ is isomorphic to $\mathfrak{L}$. The choice of the notation $d\alpha$ for this kind of contractions is justified by the fact that it is the coboundary of the map $\alpha$ if we consider $\mathbb{F}^{\times}$ as a $G$-module with trivial action. Furthermore, when $\alpha$ is a group homomorphism then $d\alpha=1$.
Graded degenerations. Let $G$ be an abelian group. Consider a $G$-grading $\Gamma$ on a finite-dimensional vector space $V$ over an algebraically closed field $\mathbb{F}$, $V=\bigoplus_{g\in G}V_{g}$, with $\dim V_{g}=n_{g}\in\mathbb{N}_{0}$. Denote $\underline{n}=(n_{g})_{g\in G}$ and let $\mathcal{L}_{\underline{n}}(\mathbb{F})$ be the variety of Lie algebra laws on $V$ that make it a $G$-graded algebra. The set $\mathcal{L}_{\underline{n}}(\mathbb{F})$ is an algebraic set in the affine space $\wedge^{2}V^{*}\otimes V$, which is invariant under the natural action of the subgroup $\prod_{g\in G}{\rm GL}(V_{g})$ of ${\rm GL}(V)$. We will consider the subgroup $D(\Gamma)$ of $\prod_{g\in G}{\rm GL}(V_{g})$ consisting of the operators that act as (nonzero) scalars on each $V_{g}$. It is clear that $D(\Gamma)$ is an algebraic torus of dimension $|{\rm Supp}(\Gamma)|$.
We can define graded degenerations in a way similar to how ordinary degenerations were defined in Section 2.1, just replacing ${\rm GL}(V)$ with the group $D(\Gamma)$. It turns out that the graded degenerations $\mu_{0}$ of a Lie algebra law $\mu\in\mathcal{L}_{\bar{n}}(\mathbb{F})$ depend on the set $S={\rm Supp}_{2}(\Gamma)$ defined by
${\rm Supp}_{2}(\Gamma):=\{(g,h)\in G\times G\mid\mu(V_{g},V_{h})\neq 0\},$
which we will call the second support of the grading $\Gamma$. Note that, in the case of Lie algebras, the order of $g$ and $h$ does not matter, so ${\rm Supp}_{2}(\Gamma)$ can be regarded as a subset of the set $\mathcal{P}_{G}$ of unordered pairs $\{g,h\}$. We discuss this in detail in Section 4.
Proposition 7.
Any subset of $\mathcal{P}_{G}$ is the second support of a $G$-graded Lie algebra. Moreover, if $G$ is finite, this Lie algebra can be chosen finite-dimensional.
Proof.
Consider the free Lie algebra $\mathfrak{F}$ on the generators $x_{g}$ and $y_{g}$, where $g$ runs through $G$. We can define a $G$-grading on $\mathfrak{F}$ by setting the $G$-degree of generators $x_{g}$ and $y_{g}$ to be $g$. Given a set $S\subset\mathcal{P}_{G}$, consider the ideal $I_{S}$ of $\mathfrak{F}$ generated by $\mathfrak{F}^{3}$ (the third term of the lower central series of $\mathfrak{F}$) and the commutators $[\mathfrak{F}_{g},\mathfrak{F}_{h}]$ for all $\{g,h\}\notin S$. Then the quotient $\mathfrak{L}:=\mathfrak{F}/I_{S}$ is a $G$-graded Lie algebra, whose second support is $S$. Moreover, if $G$ is finite, the algebra $\mathfrak{L}$ is finite-dimensional. ∎
Remark 8.
Replacing ${\rm GL}(V)$ with the product $\prod_{g\in G}{\rm GL}(V_{g})$, one can define the concept of degeneration of a $G$-graded Lie algebra, which is different from graded degeneration defined above.
Equivalences of graded contractions. For a specific fixed grading $\Gamma$ of a Lie algebra $\mathfrak{L}$, graded contractions are studied modulo several natural equivalences, see [9, Definitions 2.4, 2.6, 2.11]. We translate these notions into the generic case.
Definition 9.
Two graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ are called equivalent, written $\varepsilon\sim\varepsilon^{\prime}$, if, for any $G$-graded Lie algebra $\mathfrak{L}$, there exists an equivalence of graded Lie algebras $\varphi\colon\mathfrak{L}^{\varepsilon}\to\mathfrak{L}^{\varepsilon^{\prime}}$, i.e., a Lie algebra isomorphism mapping each homogeneous component of $\mathfrak{L}^{\varepsilon}$ onto a homogeneous component of $\mathfrak{L}^{\varepsilon^{\prime}}$: for any $g\in G$ there is $h\in G$ such that $\varphi\left(\mathfrak{L}^{\varepsilon}_{g}\right)=\mathfrak{L}^{\varepsilon^{\prime}}_{h}$.
Definition 10.
Two contractions $\varepsilon$ and $\varepsilon^{\prime}$ are called strongly equivalent, written $\varepsilon\approx\varepsilon^{\prime}$, if, for any $G$-graded Lie algebra $\mathfrak{L}$, there exists an isomorphism of $G$-graded Lie algebras (for short, a graded isomorphism) $\varphi\colon\mathfrak{L}^{\varepsilon}\to\mathfrak{L}^{\varepsilon^{\prime}}$, i.e., an isomorphism of Lie algebras mapping each homogeneous component of $\mathfrak{L}^{\varepsilon}$ onto the homogeneous component of $\mathfrak{L}^{\varepsilon^{\prime}}$ of the same degree: $\varphi\big{(}\mathfrak{L}^{\varepsilon}_{g}\big{)}=\mathfrak{L}^{\varepsilon^{\prime}}_{g}$ for any $g\in G$.
It is clear that any two strongly equivalent contractions are equivalent, but not conversely, see [9, Example 2.5].
For arbitrary two graded contractions, $\varepsilon$ and $\varepsilon^{\prime}$, one can define their product $\varepsilon\varepsilon^{\prime}\colon G\times G\to\mathbb{F}$ to be simply the point-wise product of functions,
$(\varepsilon\varepsilon^{\prime})(g,h):=\varepsilon(g,h)\varepsilon^{\prime}(g,h).$
It is clear that $\varepsilon\varepsilon^{\prime}$ is also a graded contraction. In particular, if $\varepsilon$ is a graded contraction, then $\varepsilon\,d\alpha$ is a graded contraction for any function $\alpha\colon G\to\mathbb{F}^{\times}$. This gives rise to the third notion of equivalence of graded contractions.
Definition 11.
Two graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ are called equivalent via normalization, written $\varepsilon\sim_{n}\varepsilon^{\prime}$, if there exists a map $\alpha\colon G\to\mathbb{F}^{\times}$ such that $\varepsilon^{\prime}=\varepsilon\,d\alpha$.
Definition 11 can be rephrased as follows: two maps $\varepsilon,\varepsilon^{\prime}\colon G\times G\to\mathbb{F}$ satisfying the “two-cocycle” condition $(ii^{\prime})$ of Proposition 5 are equivalent via normalization if and only if they are cohomologous. This provides us with the cohomological framework for the study of graded contractions, which is discussed in more detail in Section 3.
The equivalence via normalization implies the strong equivalence, $\varepsilon\sim_{n}\varepsilon^{\prime}\Rightarrow\varepsilon\approx\varepsilon^{\prime}$. The graded isomorphism $\varphi\colon\mathfrak{L}^{\varepsilon}\to\mathfrak{L}^{\varepsilon^{\prime}}$ is given by
$\varphi(x)=\alpha(g)x\quad\forall x\in\mathfrak{L}_{g}$
for all $G$-graded Lie algebras $\mathfrak{L}$. The converse implication, $\varepsilon\approx\varepsilon^{\prime}\Rightarrow\varepsilon\sim_{n}\varepsilon^{\prime}$, is conjectured in [39, Conjecture 2.15]. While $\mathfrak{L}^{\varepsilon}$ may be isomorphic to $\mathfrak{L}^{\varepsilon^{\prime}}$, for specific $\mathfrak{L}$, without $\varepsilon$ and $\varepsilon^{\prime}$ being equivalent via normalization (see, for example, [9, Example 2.12]), we are not aware of counterexamples to this conjecture. Moreover, we show that a version of this conjecture holds when the graded isomorphism for all $\mathfrak{L}$ is replaced by its functorial counterpart, i.e., natural graded isomorphism. This is discussed in detail in Section 5.
3 Graded contractions via group cohomology
In [28, Section 3] Moody and Patera initiated the study of “generic” $G$-graded Lie algebra contractions, which they defined as functions $\varepsilon\colon G\times G\to\mathbb{F}$ satisfying conditions $(i^{\prime})$ and $(ii^{\prime})$ of Proposition 5, from the perspective of semigroup cohomology. Indeed, since the field $\mathbb{F}$ is a semigroup with respect to multiplication, it can be regarded as a $G$-semimodule with trivial action. In view of conditions $(i^{\prime})$ and $(ii^{\prime})$, a contraction $\varepsilon$ is a symmetric two-cocycle with values in this $G$-semimodule, moreover the equivalence via normalization coincides with the relation of “being cohomologous”. Since cohomology theory is more developed for groups than semigroups, we will follow a group-cohomology approach, which was inspired by the work of Weimar–Woods [39], even though group cohomology does not appear there explicitly.
Definition 12.
Let $G$ be an abelian group. We denote by $\mathcal{P}_{G}$ the set of unordered pairs of elements of $G$,
$\mathcal{P}_{G}:=\{\{g,h\}\subset G\}.$
If the group $G$ is finite, the cardinality of $\mathcal{P}_{G}$ equals $|G|(|G|+1)/2$. In view of condition $(i^{\prime})$ of Proposition 5, a $G$-graded contraction $\varepsilon$ can be considered as a function $\mathcal{P}_{G}\to\mathbb{F}$. In particular, the support of $\varepsilon$ is the set
$S(\varepsilon):=\{\{g,h\}\in\mathcal{P}_{G}\mid\varepsilon(g,h)\neq 0\}.$
The contraction $\varepsilon$ is said to be without zeros if $S(\varepsilon)=\mathcal{P}_{G}$, and with zeros otherwise.
Conditions $(i^{\prime})$ and $(ii^{\prime})$ tell us that a set $S\subset\mathcal{P}_{G}$ is a support of a contraction if and only if $\{g,h\},\{gh,k\}\in S$ implies $\{h,k\},\{g,hk\}\in S$ for all $g,h,k\in G$. We denote the set of all such subsets of $\mathcal{P}_{G}$ by $\mathcal{S}(G)$.
Finding supports is an essential step in the classification of $G$-graded contractions. While testing if a given subset $S$ of $\mathcal{P}_{G}$ belongs to $\mathcal{S}(G)$ is straightforward for finite $G$, finding all elements of the set $\mathcal{S}(G)$ is a challenging computational problem because of the growth rate of the size of the power set of $\mathcal{P}_{G}$ as a function of $N:=|G|$.
To find the set $\mathcal{S}(G)$, we develop an efficient approach based on some elementary facts from commutative algebra and affine algebraic geometry. It relies on the efficient algorithms of computing Gröbner bases and taking quotient rings in the computer algebra system Singular [7]. We discuss this algorithm in detail in Section 4.
Remark 13.
The set $\mathcal{S}(G)$ is closed with respect to taking intersections, i.e., if $S_{1},S_{2}\in\mathcal{S}(G)$, then the intersection $S_{1}\cap S_{2}$ also belongs to $\mathcal{S}(G)$. Thus, $\mathcal{S}(G)$ is a meet semilattice with respect to inclusion. This observation suggests addressing the problem from the perspective of the semilattice structure of $\mathcal{S}(G)$. Indeed, given all elements of $\mathcal{S}(G)$ that are maximal with the property of not containing a fixed element of $\mathcal{P}_{G}$, all other elements of $\mathcal{S}_{G}$ can be generated by taking all possible intersections. This approach avoids solving polynomial equations that define contractions.
Let $\mathcal{F}_{G}$ be the free abelian group, written multiplicatively, on the alphabet $\mathcal{P}_{G}$, and let $\mathcal{R}_{G}$ be the set of relators of the form
$\{g,h\}\{gh,k\}\{h,k\}^{-1}\{g,hk\}^{-1},$
for all $g,h,k\in G$. Any contraction with support $\mathcal{P}_{G}$ (i.e., without zeros) defines a unique homomorphism $\mathcal{F}_{G}\to\mathbb{F}^{\times}$. Under this homomorphism, the elements of the subgroup $\langle\mathcal{R}_{G}\rangle$ are mapped to $1\in\mathbb{F}$, so we will call them the higher-order identities. Thus, the group of contractions with support $\mathcal{P}_{G}$ may be identified with the group of multiplicative characters of the group $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$, ${\rm Hom}_{\mathbb{Z}}(\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle,\mathbb{F}^{\times})$. In what follows, we simply write ${\rm Hom}(A,B)$ for the group of abelian group homomorphisms $A\to B$.
Since we are interested in contractions with zeros, it is important to note that every relation $r=1$ for $r\in\langle\mathcal{R}_{G}\rangle$ can be uniquely rewritten in the form $r_{1}=r_{2}$, where $r_{1}$ and $r_{2}$ are monomials in two disjoint sets of generators (without using inverses). We note that the degrees of these two monomials are necessarily equal, because the homomorphism $\mathcal{F}_{G}\to\mathbb{Z}$ that maps each generator of $\mathcal{F}_{G}$ to $1$ sends all elements of $\mathcal{R}_{G}$ and hence the subgroup $\langle\mathcal{R}_{G}\rangle$ to zero.
The following list of definitions is an interpretation of definitions from [39] within the framework introduced above.
Definition 14.
For any subset $S\subset\mathcal{P}_{G}$ let $\mathcal{F}_{S}$ be the free abelian group generated by $S$.
  1. 1.
    The elements of the set $\mathcal{R}_{S}:=\mathcal{R}_{G}\cap\mathcal{F}_{S}$ are called surviving defining relations. The elements of the subgroup $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ are called surviving higher-order identities.
  2. 2.
    A subset $B\subset\mathcal{P}_{G}$ is called independent if its image under the projection $\mathcal{F}_{G}\twoheadrightarrow\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ is $\mathbb{Z}$-linearly independent.
  3. 3.
    A subset $B\subset S$ is called independent (resp. quasi-independent) with respect to $S$ if the image of $B$ under the projection $\mathcal{F}_{S}\twoheadrightarrow\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ (resp. $\mathcal{F}_{S}\twoheadrightarrow\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$) is $\mathbb{Z}$-linearly independent.
  4. 4.
    We denote by $N^{\prime}$ (resp. $N^{\prime\prime}$) the free rank of the group $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ (resp. $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$), i.e., the cardinality of a maximal subset of $S$ that is independent (resp. quasi-independent) with respect to $S$.
Remark 15.
In [39], only supports of contractions, i.e. $S\in\mathcal{S}(G)$, were considered, but the definitions make sense for any subset of $\mathcal{P}_{G}$. Also, in [39], the relations $r_{1}=r_{2}$ of degrees $1$ and $2$ were excluded from higher-order identities because, for $S\in\mathcal{S}(G)$, they belong to $\langle\mathcal{R}_{S}\rangle$, i.e., follow from surviving defining relations. Indeed, a surviving identity of degree $1$ has the form $\{g_{1},h_{1}\}=\{g_{2},h_{2}\}$, where both sides are distinct and belong to $S$. Consider the group homomorphism $f\colon\mathcal{F}_{G}\to\mathbb{Z}G$ defined by $\{g,h\}\mapsto g+h-gh$. Since $f$ sends $\mathcal{R}_{G}$ to zero, it must kill all elements of $\langle\mathcal{R}_{G}\rangle$, in particular $\{g_{1},h_{1}\}\{g_{2},h_{2}\}^{-1}$, so we have $g_{1}+h_{1}-g_{1}h_{1}=g_{2}+h_{2}-g_{2}h_{2}$. This implies that the degree $1$ identity has the form $\{e,e\}=\{e,g\}$ with both sides belonging to $S$, so it follows from the surviving defining relation $\{e,e\}\{e,g\}=\{e,g\}\{e,g\}$. Similar considerations work for degree $2$ showing that the surviving degree $2$ identities are in fact surviving defining relations.
The inclusion $\langle\mathcal{R}_{S}\rangle\hookrightarrow\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ induces a surjection $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle\twoheadrightarrow\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$. This gives us that $N^{\prime\prime}\geqslant N^{\prime}$. Since the free rank of an abelian group $A$ is the dimension of the $\mathbb{Q}$-vector space $A\otimes_{\mathbb{Z}}\mathbb{Q}$, the short exact sequence of abelian groups
$0\rightarrow(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle\rightarrow\mathcal{F}_{S}/\mathcal{\langle}\mathcal{R}_{S}\rangle\rightarrow\mathcal{F}_{S}/(\mathcal{\langle}\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})\rightarrow 0$
implies the following equality:
$\mathop{\rm rank}\big{(}\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle\big{)}=\mathop{\rm rank}\big{(}\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})\big{)}+\mathop{\rm rank}\big{(}(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle\big{)},$
hence $\mathop{\rm rank}\big{(}(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle\big{)}=N^{\prime\prime}-N^{\prime}$ if $G$ is finite.
For any abelian group $A$, viewed as a trivial $G$-module, the set of symmetric two-cocycles $Z^{2}_{sym}(G,A)$ with values in $A$ can be identified with ${\rm Hom}(\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle,A)$. It is well-known that
$Z^{2}_{sym}(G,A)/B^{2}(G,A)\simeq{\rm Ext}(G,A),$ (4)
where ${\rm Ext}(G,A)={\rm Ext}_{\mathbb{Z}}(G,A)$, which can be computed using the cochain complex ${\rm Hom}_{\mathbb{Z}}(C_{\bullet},A)$ for any free resolution $C_{\bullet}$ of the abelian group $G$.
Consider the following complex of abelian groups:
$C_{\bullet}\colon\quad C_{3}\to C_{2}\to C_{1}\to C_{0}\to 0,$
where $C_{0}:=G$, $C_{1}:=\mathbb{Z}G$, $C_{2}:=\mathcal{F}_{G}$ and $C_{3}:=\mathbb{Z}(G\times G\times G)$ and the differential $\partial_{\bullet}$ is defined as follows:
$\partial_{1}\colon C_{1}\to C_{0},\quad\sum_{g\in G}n_{g}g\mapsto\prod_{g\in G}g^{n_{g}},$
$\partial_{2}\colon C_{2}\to C_{1},\quad\{g,h\}\mapsto g+h-gh,$
$\partial_{3}\colon C_{3}\to C_{2},\quad(g,h,k)\mapsto\{h,k\}\{g,hk\}\{g,h\}^{-1}\{gh,k\}^{-1}.$
Lemma 16.
The sequence $C_{\bullet}$ is exact, i.e., a free resolution of $C_{0}=G$.
Proof.
The exactness at the terms $C_{0}$ and $C_{1}$ is clear. To show exactness at $C_{2}$, we can apply the exact faithful functor ${\rm Hom}(\cdot,\mathbb{Q}/\mathbb{Z})$ and see that the resulting sequence is exact at ${\rm Hom}(C_{2},\mathbb{Q}/\mathbb{Z})$, i.e. $Z^{2}_{sym}(G,\mathbb{Q}/\mathbb{Z})=B^{2}(G,\mathbb{Q}/\mathbb{Z})$, because, in view of (4), the quotient $Z^{2}_{sym}(G,\mathbb{Q}/\mathbb{Z})/B^{2}(G,\mathbb{Q}/\mathbb{Z})$ is ${\rm Ext}(G,\mathbb{Q}/\mathbb{Z})$, which is trivial for any $G$. ∎
Corollary 17.
$\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ is a free abelian group of rank $\leqslant|G|$. If $N:=|G|$ is finite then $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ has rank $N$.
Proof.
By Lemma 16, $\langle\mathcal{R}_{G}\rangle=\mathop{\rm im}\partial_{3}=\ker\partial_{2}$, so $\partial_{2}$ induces an embedding $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle\hookrightarrow C_{1}$. Since $C_{1}$ is a free abelian group of rank $N$, it follows that $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ is a free abelian group of rank $\leqslant N$. If $G$ is finite, then the rank of $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ is equal to $N$, since the cokernel of $\partial_{2}$ is finite (isomorphic to $G$). ∎
Now consider a subset $S\subset\mathcal{P}_{G}$ and the corresponding groups $\mathcal{F}_{S}$, $\mathcal{R}_{S}$ and $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ as in Definition 14. Recall that the free ranks of the quotient groups $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ and $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ are denoted by $N^{\prime\prime}$ and $N^{\prime}$, respectively. Since the group $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ is isomorphic to the image of $\mathcal{F}_{S}$ under the quotient map $\mathcal{F}_{G}\to\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$, it is a free abelian group of rank $N^{\prime}\leqslant N$.
Remark 18.
It is claimed in [39, Theorem 6.5] that if $S$ belongs to $\mathcal{S}(G)$ then the group $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$ is also free (of rank $Q=N^{\prime\prime}-N^{\prime}$), but the proof appears to be incorrect (in the last paragraph of the proof, a function taking values in roots of unity is treated as a constant). Since we know that $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ is free, the freeness of $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$ is equivalent to the freeness of $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$. We carried out numerical experiments for small groups $G$ and verified that the group $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ is free not only for $S\in\mathcal{S}(G)$, but for all subsets $S$ of $\mathcal{P}_{G}$.
For $S\in\mathcal{S}(G)$, the significance of the group $\mathcal{Z}_{S}:=\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ is that the group of multiplicative characters ${\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})$ is isomorphic to the group of contractions whose support is $S$. Indeed, given $\gamma\in{\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})$, we define $\varepsilon\colon G\times G\to\mathbb{F}$ by
$\varepsilon(g,h):=\begin{cases}\gamma(\{g,h\})&\mbox{if}\ \{g,h\}\in S,\\ 0&\mbox{otherwise}.\end{cases}$
Condition $(i^{\prime})$ of Proposition 5 is obviously satisfied. Relations $\mathcal{R}_{S}$ imply that condition $(ii^{\prime})$ is satisfied when both sides are nonzero (there are no “strong violations” in the terminology of [39]), while the fact that $S\in\mathcal{S}(G)$ implies that the sides of $(ii^{\prime})$ are either both zero or both nonzero (there are no “weak violations” in the terminology of [39]). Conversely, every graded contraction with support $S$ has the above form.
The group of equivalence classes of these contractions is isomorphic to the quotient by the image of $B^{2}(G,\mathbb{F}^{\times})$ under the restriction map ${\rm Fun}(\mathcal{P}_{G},\mathbb{F}^{\times})\to{\rm Fun}(S,\mathbb{F}^{\times})$, where the latter can be identified with the group ${\rm Hom}(\mathcal{F}_{S},\mathbb{F}^{\times})$ since $\mathcal{F}_{S}$ is the free abelian group generated by $S$. For any abelian group $A$, we denote by $B^{2}_{S}(A)$ the image of $B^{2}(G,A)$ in ${\rm Hom}(\mathcal{F}_{S},A)$, which lies in ${\rm Hom}(\mathcal{Z}_{S},A)$. We also denote
$H^{2}_{S}(A):={\rm Hom}(\mathcal{Z}_{S},A)/B^{2}_{S}(A).$
Then, for $S\in\mathcal{S}(G)$, the group $H^{2}_{S}(\mathbb{F}^{\times})={\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})$ classifies contractions with support $S$ up to equivalence via normalization.
We want to compare the groups $\mathcal{Z}_{S}$ and $\mathcal{Z}_{G}:=\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$. There exists a unique group homomorphism $f\colon\mathcal{Z}_{S}\to\mathcal{Z}_{G}$ making the following diagram commutative:
00{0}${\langle\mathcal{R}_{S}\rangle}$${\mathcal{F}_{S}}$${\mathcal{Z}_{S}}$00{0}00{0}${\langle\mathcal{R}_{G}\rangle}$${\mathcal{F}_{G}}$${\mathcal{Z}_{G}}$00{0}$\scriptstyle{f}$
Denote $\mathcal{K}_{S}:=\ker f$, $\mathcal{I}_{S}:=\mathop{\rm im}f$ and $\mathcal{C}_{S}:=\mathop{\rm coker}f$. Then
$\mathcal{K}_{S}=(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle,$
$\mathcal{I}_{S}=(\langle\mathcal{R}_{G}\rangle\mathcal{F}_{S})/\langle\mathcal{R}_{G}\rangle\simeq\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}),$
$\mathcal{C}_{S}\simeq\mathcal{F}_{G}/(\langle\mathcal{R}_{G}\rangle\mathcal{F}_{S}).$
Recall that when the group $G$ has finite order $N$, the group $\mathcal{F}_{G}$ is free abelian of rank $N(N+1)/2$. The groups $\mathcal{F}_{S}$, $\langle\mathcal{R}_{G}\rangle$ and $\langle\mathcal{R}_{S}\rangle$ are subgroups of $\mathcal{F}_{G}$ defined by explicit sets of generators. Thus the groups $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$, $\mathcal{K}_{S}$, $\mathcal{I}_{S}$ and $\mathcal{C}_{S}$ can be computed using the Smith canonical form of an integer matrix.
Theorem 19.
For any $S\subset\mathcal{P}_{G}$ and any abelian group $A$, we have two short exact sequences:
$0\longrightarrow$ $\frac{{\rm Hom}(\mathcal{I}_{S},A)}{B^{2}_{S}(A)}\longrightarrow H^{2}_{S}(A)\longrightarrow{\rm Hom}(\mathcal{K}_{S},A)\longrightarrow 0,$ (5)
${\rm Ext}(G,A)\longrightarrow\frac{{\rm Hom}(\mathcal{I}_{S},A)}{B^{2}_{S}(A)}\longrightarrow{\rm Ext}(\mathcal{C}_{S},A)\longrightarrow 0,$ (6)
where the kernel of the map ${\rm Ext}(G,A)\to{\rm Hom}(\mathcal{I}_{S},A)/B^{2}_{S}(A)$ consists of the classes of symmetric two-cocycles whose restriction to $S$ is in $B^{2}_{S}(A)$. Moreover, (5) is split.
Proof.
Consider the short exact sequence $0\to\mathcal{K}_{S}\to\mathcal{Z}_{S}\;{\buildrel f\over{\to}}\;\mathcal{I}_{S}\to 0$. It splits since $\mathcal{I}_{S}$ is free, being a subgroup of $\mathcal{Z}_{G}$, which is a free abelian group by Corollary 17. This gives the following split exact sequence:
$0\to{\rm Hom}(\mathcal{I}_{S},A)\to{\rm Hom}(\mathcal{Z}_{S},A)\to{\rm Hom}(\mathcal{K}_{S},A)\to 0$ (7)
Since ${\rm Ext}(\mathcal{Z}_{G},A)=0$, the short exact sequence $0\to\mathcal{I}_{S}\to\mathcal{Z}_{G}\to\mathcal{C}_{S}\to 0$ gives
$0\to{\rm Hom}(\mathcal{C}_{S},A)\to{\rm Hom}(\mathcal{Z}_{G},A)\to{\rm Hom}(\mathcal{I}_{S},A)\to{\rm Ext}(\mathcal{C}_{S},A)\to 0$ (8)
The composition of ${\rm Hom}(\mathcal{Z}_{G},A)\to{\rm Hom}(\mathcal{I}_{S},A)$ from (8) and ${\rm Hom}(\mathcal{I}_{S},A)\to{\rm Hom}(\mathcal{Z}_{S},A)$ from (7) is the map $f^{*}$, which corresponds to the restriction of symmetric two-cocycles from $\mathcal{P}_{G}$ to $S$. By definition, it maps $B^{2}(G,A)$ onto $B^{2}_{S}(A)$. Recall that ${\rm Hom}(\mathcal{Z}_{G},A)/B^{2}(G,A)\simeq{\rm Ext}(G,A)$, so sequence (8) gives us (6) . Since $H^{2}_{S}(A)={\rm Hom}(\mathcal{Z}_{S},A)/B^{2}_{S}(A)$, sequence (7) gives us (5). ∎
Note that for contractions without zeros, Theorem 19 implies that $H^{2}_{S}(A)={\rm Ext}(G,A)$ as expected, since in this case $S=\mathcal{P}_{G}$ and therefore $f$ is the identity map.
Remark 20.
In the terminology of [39], the subgroup ${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})$ in sequence (5) corresponds to the classes of contractions with support $S$ that satisfy all surviving higher-order identities, i.e., take value $1$ when applied to the elements of the subgroup $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ of $\mathcal{F}_{S}$.
The following corollary was obtained in the case $\mathbb{F}=\mathbb{C}$ in [39, Theorem 7.1].
Corollary 21.
If $\mathbb{F}$ is algebraically closed then $H^{2}_{S}(\mathbb{F}^{\times})\simeq{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times})$.
Proof.
Since $\mathbb{F}^{\times}$ is a divisible group, both ${\rm Ext}(G,\mathbb{F}^{\times})$ and ${\rm Ext}(\mathcal{C}_{S},\mathbb{F}^{\times})$ are trivial, so sequence (6) implies ${\rm Hom}(\mathcal{I}_{S},A)/B^{2}_{S}(A)=0$. Consequently, sequence (5) gives the result. ∎
To treat the real case, we recall some standard results from homological algebra. For any abelian group $H$ (which we write multiplicatively), we use the following notation:
$H_{[n]}:=\{h\in H\mid h^{n}=e\}\quad\mbox{and}\quad H^{[n]}:=\{h^{n}\mid h\in H\}.$
Then ${\rm Ext}(H,\mathbb{Z}/n\mathbb{Z})$ is isomorphic to the dual group, ${\rm Hom}(H_{[n]},\mathbb{Q}/\mathbb{Z})$, of the $n$-periodic part of $H$. This can be seen, for example, from the injective resolution
$\mathbb{Z}/n\mathbb{Z}\to\mathbb{Q}/\mathbb{Z}\to\mathbb{Q}/\mathbb{Z}\to 1,$
where the map $\mathbb{Q}/\mathbb{Z}\to\mathbb{Q}/\mathbb{Z}$ is $x\mapsto nx$ (writing additively). It gives that ${\rm Ext}(H,\mathbb{Z}/n\mathbb{Z})$ is isomorphic to the quotient ${\rm Hom}(H,\mathbb{Q}/\mathbb{Z})/{\rm Hom}(H,\mathbb{Q}/\mathbb{Z})^{[n]}$, which is in turn isomorphic to ${\rm Hom}(H_{[n]},\mathbb{Q}/\mathbb{Z})$, by applying the exact functor ${\rm Hom}(\cdot,\mathbb{Q}/\mathbb{Z})$ to the exact sequence
$1\to H_{[n]}\to H\to H.$
If $\mathbb{F}$ is a real closed field (for example, $\mathbb{R}$), then the multiplicative group of $\mathbb{F}$ is the direct sum of {±$\{\pm 1\}\simeq\mathbb{Z}/2\mathbb{Z}$ and the divisible group of positive elements $\mathbb{F}_{>0}$, hence the group ${\rm Ext}(H,\mathbb{F}^{\times})$ is isomorphic to the dual group of $H_{[2]}$. Then sequence (6) takes the form
${\rm Hom}(G_{[2]},\{\pm 1\})\longrightarrow\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})}{B^{2}_{S}(\mathbb{F}^{\times})}\longrightarrow{\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})\longrightarrow 0.$ (9)
Corollary 22.
If $\mathbb{F}$ is real closed then
$H^{2}_{S}(\mathbb{F}^{\times})\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{B^{2}_{S}(\{\pm 1\})}\oplus{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times}),$ (10)
where ${{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}/{B^{2}_{S}(\{\pm 1\})}$ is an elementary abelian $2$-group. If $G$ is finite, then the rank of this group (i.e., dimension as a vector space over the field of $2$ elements) is at most $\mathop{\rm rank}G_{[2]}+N^{\prime}$.
Proof.
Since $G_{[2]}$ and $(\mathcal{C}_{S})_{[2]}$ are elementary $2$-groups, (9) implies that ${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})$ has period at most $4$. On the other hand, using the direct sum decomposition of $\mathbb{F}^{\times}$, we get that
$\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})}{B^{2}_{S}(\mathbb{F}^{\times})}\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{B^{2}_{S}(\{\pm 1\})}\oplus\frac{{\rm Hom}(\mathcal{I}_{S},\mathbb{F}_{>0})}{B^{2}_{S}(\mathbb{F}_{>0})}.$
Since $\mathcal{I}_{S}$ is a free abelian group, the second summand above is a divisible group and hence cannot have finite period unless it is trivial. Therefore,
${\rm Hom}(\mathcal{I}_{S},\mathbb{F}^{\times})/B^{2}_{S}(\mathbb{F}^{\times})\simeq\frac{{\rm Hom}(\mathcal{I}_{S},\{\pm 1\})}{B^{2}_{S}(\{\pm 1\})}$
is an elementary 2-group. If $N:=|G|$ is finite then, in view of sequence (9), the rank of this group is at most $\mathop{\rm rank}G_{[2]}+N^{\prime}$, because $\mathcal{C}_{S}$ is isomorphic to the quotient of the free abelian group $\mathcal{Z}_{G}=\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ of rank $N$ by its subgroup $\mathcal{I}_{S}\simeq\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})$ of rank $N^{\prime}$. The split sequence (5) completes the proof. ∎
Remark 23.
Recall from Remark 20 that, in the terminology of [39], the first summand in equation (10) corresponds to the classes of contractions with support $S$ that satisfy all surviving higher-order identities. Such classes are described, for $\mathbb{F}=\mathbb{R}$, in [39, Theorem 7.1] in terms of “sign invariants”. In sequence (9), the quotient ${\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})$ correspond to the “sign invariants of the first kind” as follows. Fix a set-theoretic section $\xi$ of the projection $\mathcal{Z}_{G}\to\mathcal{C}_{S}$. Given a contraction $\varepsilon$ that satisfies all surviving higher-order identities, i.e., $\varepsilon\in{\rm Hom}(\mathcal{I}_{S},\mathbb{R}^{\times})$, we define the two-cocycle $\alpha_{\varepsilon}\in Z^{2}_{sym}(\mathcal{C}_{S},\{\pm 1\})$ by
$\alpha_{\varepsilon}(c_{1},c_{2}):=\mathop{\rm sgn}\nolimits\big{(}\varepsilon(\xi(c_{1})\xi(c_{2})\xi(c_{1}c_{2})^{-1})\big{)},\quad c_{1},c_{2}\in\mathcal{C}_{S},$
where $\mathop{\rm sgn}\nolimits$ denotes the usual sign function, which is the projection ×{±$\mathbb{R}^{\times}\twoheadrightarrow\{\pm 1\}$. The class $[\alpha_{\varepsilon}]\in{\rm Ext}(\mathcal{C}_{S},\{\pm 1\})$ is the image of $[\varepsilon]$ under the homomorphism ${\rm Hom}(\mathcal{I}_{S},\mathbb{R}^{\times})/B^{2}_{S}(\mathbb{R}^{\times})\to{\rm Ext}(\mathcal{C}_{S},\{\pm 1\})$ from sequence (6). The corresponding element $\delta_{\varepsilon}\in{\rm Hom}((\mathcal{C}_{S})_{[2]},\{\pm 1\})$ is given by
$\delta_{\varepsilon}(c):=\alpha_{\varepsilon}(c,c)\alpha_{\varepsilon}(e,e)=\mathop{\rm sgn}\nolimits\big{(}\varepsilon(\xi(c)^{2})\big{)},\quad c\in\mathcal{(}\mathcal{C}_{S})_{[2]}.$
The values of $\delta_{\varepsilon}$ are the “sign invariants of the first kind” assigned to $\varepsilon$.
4 Variety of graded contractions
Now we approach $G$-graded contractions from the perspective of commutative algebra and affine algebraic geometry. More specifically, we consider condition $(ii^{\prime})$ of Proposition 5 as a relation in the ring of polynomials $\mathbb{F}[x_{i}\mid i\in\mathcal{P}_{G}]$, whose variables are labeled, in view of condition $(i^{\prime})$, by the set $\mathcal{P}_{G}$ of unordered pairs $\{g,h\}\subset G$. To avoid technical issues, the ground field $\mathbb{F}$ in this section is assumed to be algebraically closed.
Cellular decomposition of the variety of graded contractions. Let $G$ be a finite abelian group and consider the affine space $\mathbb{A}_{G}$, whose coordinates are indexed by the set $\mathcal{P}_{G}$, so the coordinate algebra is $\mathbb{F}[\mathbb{A}_{G}]=\mathbb{F}[x_{i}\mid i\in\mathcal{P}_{G}]$. Then the set $X$ of all $G$-graded contractions can be regarded as a subset of $\mathbb{A}_{G}$. Moreover, $X$ is an affine algebraic variety determined by the equations
$x_{\{g,h\}}x_{\{g,hk\}}-x_{\{h,k\}}x_{\{g,hk\}}=0\quad\mbox{for all}\quad g,h,k\in G.$ (11)
Let $I\subset\mathbb{F}[\mathbb{A}_{G}]$ be the ideal generated by the left-hand sides of the above equations. This ideal is binomial, meaning it is generated by polynomials with two terms. Binomial ideals have a number of remarkable properties, for instance, the radical and the associated primes of a binomial ideal are again binomial. Eisenbud and Sturmfels [12] provide a detailed study of these and other properties and characterize algebraic sets defined by binomial ideals in terms of their cellular decomposition, see [12, Theorem 4.2]. We will subsequently study the cellular decomposition of the variety $X$ of graded contractions. It should be noted that the ideal $I$ is usually not radical, which can be verified using Singular computer algebra system [7]. For instance, $I$ is radical when $G$ is of order $2$, $3$, $4$ or $5$, but not $6$.
For any subset $S\subset\mathcal{P}_{G}$, we consider the affine subspace $\mathbb{A}_{S}:=\{x\in\mathbb{A}_{G}\colon x_{i}\neq 0~{}\forall i\in S\}\subset\mathbb{A}_{G}$ and the torus ${\rm T}_{S}:=\{x\in\mathbb{A}_{S}\colon x_{i}\neq 0~{}\forall i\in S\}$, which is a principal open set in $\mathbb{A}_{S}$. Since $x\in{\rm T}_{S}$ if and only if $\mathop{\rm supp}x=S$, where $\mathop{\rm supp}x=\{i\in\mathcal{P}_{G}\colon x_{i}\neq 0\}$, we have a partition $\mathbb{A}_{S}=\bigcup_{S^{\prime}\subset S}{\rm T}_{S^{\prime}}$. The tori ${\rm T}_{S}$ are called coordinate cells in [12]. The intersection $X\cap{\rm T}_{S}$ consists precisely of the graded contractions, whose support is $S$; it is nonempty if and only if $S\in\mathcal{S}(G)$. It follows that we have the following partition of $X$:
$X=\bigcup_{S\in\mathcal{S}(G)}X_{S},\quad X_{S}:=X\cap{\rm T}_{S}.$
In particular, $X_{\mathcal{P}_{G}}=X\cap{\rm T}_{\mathcal{P}_{G}}=Z^{2}_{sym}(G,\mathbb{F}^{\times})$ is a nonempty principal open set in $X$. Consider the torus ${\rm T}_{G}:={\rm Fun}(G,\mathbb{F}^{\times})$ and its action on the space $\mathbb{A}_{G}$ defined by
$\alpha\cdot(x_{\{g,h\}})_{\{g,h\}\in\mathcal{P}_{G}}=\big{(}(d\alpha)(g,h)\,x_{\{g,h\}}\big{)}_{\{g,h\}\in\mathcal{P}_{G}},$
where recall that $(d\alpha)(g,h)=\alpha(g)\alpha(h)/\alpha(gh)$. In other words, ${\rm T}_{G}$ acts through the homomorphism $d\colon{\rm T}_{G}\to{\rm T}_{\mathcal{P}_{G}}$, $\alpha\mapsto d\alpha$. This action leaves $X$ and each cell $X_{S}$ invariant. Note that the ${\rm T}_{G}$-orbit of the point $\bar{1}:=(1,\dots,1)$ is $B^{2}(G,\mathbb{F}^{\times})=Z^{2}_{sym}(G,\mathbb{F}^{\times})=X_{\mathcal{P}(G)}$.
Proposition 24.
The closure of the ${\rm T}_{G}$-orbit of $\bar{1}$, which we denote by $X(\bar{1})$, is the unique irreducible component of $X$ containing $\bar{1}$. Moreover, the points of the orbit are regular in $X(\bar{1})$.
Proof.
Since ${\rm T}_{G}$ is irreducible, so is $X(\bar{1})$, hence it is contained in an irreducible component of $X$. But it contains an open neighbourhood of $\bar{1}$ in $X$, so it must contain any irreducible component of $X$ passing through $\bar{1}$.
The second assertion is a general property of orbit closures, see discussion in Section 2.1. ∎
Example 25.
Consider the grading group $G=\mathbb{Z}_{2}$. The affine space $A_{\mathbb{Z}_{2}}$ is coordinatized by the tuple $(x_{\{0,0\}},x_{\{1,1\}},x_{\{0,1\}})$ and its coordinate algebra is $\mathbb{C}[x_{\{0,0\}},x_{\{1,1\}},x_{\{0,1\}}]$. The variety of contractions is the zero locus of the polynomials
$x_{\{0,0\}}x_{\{0,1\}}-x_{\{0,1\}}^{2},\quad x_{\{0,0\}}x_{\{1,1\}}-x_{\{1,1\}}x_{\{0,1\}}.$
The ideal $I$ generated by this polynomial is actually radical. The primary decomposition of $I$ is given by
$I=\langle x_{\{0,0\}}-x_{\{0,1\}}\rangle\cap\langle x_{\{1,1\}},x_{\{1,0\}}\rangle.$
The variety $X=V(I)$ is the union $V(x_{\{0,0\}}-x_{\{0,1\}})\cup V(x_{\{1,1\}},x_{\{1,0\}})$ of the plane $x_{\{0,0\}}-x_{\{0,1\}}=0$ and the line $x_{\{1,1\}}=x_{\{1,0\}}=0$. The supports of $\mathbb{Z}_{2}$-graded contractions are the sets
$S_{1}:=\varnothing,\quad S_{2}:=\{x_{\{0,0\}}\},\quad S_{3}:=\{x_{\{1,1\}}\},$
$S_{4}:=\{x_{\{0,0\}},x_{\{0,1\}}\},\quad S_{5}:=\{x_{\{0,0\}},x_{\{1,1\}},x_{\{0,1\}}\}.$
The cellular decomposition of $X$ is given by
$X\cap{\rm T}_{S_{1}}=(0,0,0),\quad X\cap{\rm T}_{S_{2}}=\big{(}V(x_{\{0,0\}}-x_{\{0,1\}})\cup V(x_{\{1,1\}},x_{\{1,0\}})\big{)}\cap{\rm T}_{S_{2}}$
$X\cap{\rm T}_{S_{i}}=V(x_{\{0,0\}}-x_{\{0,1\}})\cap{\rm T}_{S_{i}},\quad i=3,4,5.$
The ${\rm T}_{\mathbb{Z}_{2}}$-orbit of the point $(1,1,1)$, which corresponds to the identity function $\bar{1}$, is the cell $X\cap{\rm T}_{S_{5}}$ of $X$. Its closure is the plane $V(x_{\{0,0\}}-x_{\{0,1\}})$.
Example 26.
The computations for the grading group $G=\mathbb{Z}_{3}$ are more complicated than those presented in Example 25. The coordinate algebra is the polynomial algebra in six variables, $\mathbb{C}[x_{\{0,0\}},x_{\{1,1\}},x_{\{2,2\}},x_{\{0,1\}},x_{\{0,2\}},x_{\{1,2\}}]$. The variety $X$ of contractions is the zero locus of the polynomials
$x_{\{0,0\}}x_{\{0,1\}}-x_{\{0,1\}}^{2},\quad x_{\{0,0\}}x_{\{0,2\}}-x_{\{0,2\}}^{2},\quad x_{\{0,0\}}x_{\{1,2\}}-x_{\{0,1\}}x_{\{1,2\}},$
$x_{\{0,0\}}x_{\{1,2\}}-x_{\{0,2\}}x_{\{1,2\}},\quad x_{\{1,1\}}x_{\{0,1\}}-x_{\{1,1\}}x_{\{0,2\}},\quad x_{\{2,2\}}x_{\{0,1\}}-x_{\{2,2\}}x_{\{0,2\}},$
$x_{\{0,1\}}x_{\{1,2\}}-x_{\{0,2\}}x_{\{1,2\}},\quad x_{\{1,1\}}x_{\{2,2\}}-x_{\{0,1\}}x_{\{1,2\}},\quad x_{\{1,1\}}x_{\{2,2\}}-x_{\{0,2\}}x_{\{1,2\}}.$
The primary decomposition of $I$ is given by
$I=\langle x_{\{0,1\}}-x_{\{0,2\}},\,x_{\{1,1\}}x_{\{2,2\}}-x_{\{0,2\}}x_{\{1,2\}},\,x_{\{0,0\}}-x_{\{0,2\}}\rangle$
$\qquad\cap\langle x_{\{1,2\}},x_{\{0,2\}},x_{\{0,1\}},x_{\{2,2\}}\rangle\cap\langle x_{\{1,2\}},x_{\{0,2\}},x_{\{0,1\}},x_{\{1,1\}}\rangle$
$\qquad\cap\langle x_{\{1,2\}},x_{\{0,1\}},x_{\{2,2\}},x_{\{1,1\}},x_{\{0,0\}}-x_{\{0,2\}}\rangle$
$\qquad\cap\langle x_{\{1,2\}},x_{\{0,2\}},x_{\{2,2\}},x_{\{1,1\}},x_{\{0,0\}}-x_{\{0,1\}}\rangle.$
The irreducible components of the variety $X$ are
  • two lines: $V(x_{\{1,2\}},\,x_{\{0,2\}},\,x_{\{2,2\}},\,x_{\{1,1\}},\,x_{\{0,0\}}-x_{\{0,1\}})$ and $V(x_{\{1,2\}},\,x_{\{0,1\}},\,x_{\{2,2\}},\,x_{\{1,1\}},\newline x_{\{0,0\}}-x_{\{0,2\}})$,
  • two planes: $V(x_{\{1,2\}},x_{\{0,2\}},x_{\{0,1\}},x_{\{2,2\}})$ and $V(x_{\{1,2\}},x_{\{0,2\}},x_{\{0,1\}},x_{\{2,2\}})$,
  • dimension-three variety: $V(x_{\{0,1\}}-x_{\{0,2\}},\,x_{\{1,1\}}x_{\{2,2\}}-x_{\{0,2\}}x_{\{1,2\}},\,x_{\{0,0\}}-x_{\{0,2\}})$.
Finding supports. We now describe an algorithm to find the elements of $\mathcal{S}(G)$ and count their number. Let $J$ be the ideal in the polynomial ring $\mathbb{C}[x_{\{g,h\}}\mid g,h\in G]=\mathbb{C}[\mathbb{A}_{G}]$ generated by the binomials
$x_{\{g,h\}}x_{\{g,hk\}}-x_{\{h,k\}}x_{\{g,hk\}},\quad x_{\{g,h\}}(x_{\{g,h\}}-1)\quad\mbox{for all}\quad g,h,k\in G.$
The affine variety $V(J)$ is a finite set of points in the affine space $\mathbb{A}_{G}$. These points satisfy the determining equations for $G$-graded contractions and their coordinates are zero or one, so they correspond to the elements of $\mathcal{S}(G)$.
Now consider the algebra $\mathbb{C}[\mathbb{A}_{G}]/\sqrt{J}$ of regular functions of $V(J)$. Since $V(J)$ is merely a finite set of points, the algebra $\mathbb{C}[\mathbb{A}_{G}]/\sqrt{J}$ is finite-dimensional and its dimension is precisely the number of points in $V(J)$.
All the above computations can be efficiently carried out in Singular. Even for the order-8 grading groups, i.e., the number of varibles of $\mathbb{C}[\mathbb{A}_{G}]$ is $36$, these computations are performed within seconds. In Table 1 we present the numerical results for the number of supports depending on the grading group $G$. We also observe an interesting phenomenon that only a small part of subsets of $\mathcal{P}_{G}$ are actually supports.
Table 1: Number of supports
$G$ $\mathbb{Z}_{2}$ $\mathbb{Z}_{3}$ $\mathbb{Z}_{2}^{2}$ $\mathbb{Z}_{4}$ $\mathbb{Z}_{5}$ $\mathbb{Z}_{6}$ $\mathbb{Z}_{7}$ $\mathbb{Z}_{2}^{3}$ $\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ $\mathbb{Z}_{8}$
$|\mathcal{S}(G)|$ 5 15 $41$ $47$ $185$ $652$ $2997$ $16147$ $14179$ $13883$
$2^{|\mathcal{P}(G)|}$ 8 $2^{6}$ $2^{10}$ $2^{10}$ $2^{15}$ $2^{21}$ $2^{28}$ $2^{36}$ $2^{36}$ $2^{36}$
[…]
Orbit closures and higher-order identities. The elements of $X(\bar{1})$ are precisely the continuous contractions as defined in [28, Section 4]. In the case $\mathbb{F}=\mathbb{C}$, $\mathfrak{L}^{\varepsilon}$ is a graded degeneration of a Lie algebra $\mathfrak{L}$ for any $\varepsilon\in X(\bar{1})$, so, in particular, it is a contraction of $\mathfrak{L}$ in the sense of Definition 1. Recall from Section 3 that every relation $r=1$ for $r\in\langle\mathcal{R}_{G}\rangle$ can be uniquely rewritten in the form $r_{1}=r_{2}$, where $r_{1}$ and $r_{2}$ are monomials of the same degree in two disjoint sets of variables. Following [38, 39], we call these relations higher-order identities; they include the defining relations (11). Since the elements of $Z^{2}_{sym}(G,\mathbb{F}^{\times})$ satisfy all higher-order identities, so do the elements of the closure, $X(\bar{1})$. It follows from Theorem 19 and [38, Theorem VI.2] that any element of $X(\bar{1})$ is equivalent via normalization to a generalized Inönü–Wigner contraction.
Now let $S$ be a subset of $\mathcal{P}_{G}$ and $\bar{1}_{S}$ be its indicator function considered as the $|\mathcal{P}_{G}|$-tuple:
$(\bar{1}_{S})_{i}:=\begin{cases}1&\mbox{if }i\in S,\\ 0&\mbox{otherwise}.\end{cases}$
Note that $S\in\mathcal{S}(G)$ if and only if $\bar{1}_{S}$ belongs to $X$. In this case denote by $X(\bar{1}_{S})\subset X$ the closure of the ${\rm T}_{G}$-orbit of $\bar{1}_{S}$. In particular $S=\mathcal{P}_{G}$ gives $X(\bar{1})$. More generally, for any $\varepsilon\in X_{S}$, denote by $X(\varepsilon)$ the closure of ${\rm T}_{G}$-orbit of $\varepsilon$. Note that $X(\varepsilon)\subset\bigcup_{S^{\prime}\subset S}X_{S^{\prime}}$. The following result describes the vanishing ideal of $X(\varepsilon)$ in terms of the surviving higher-order identities, i.e., those of the form $r_{1}=r_{2}$ where both sides involve variables indexed only by $i\in S$.
Theorem 27.
Let $G$ be an abelian group and $S$ be a finite subset of $\mathcal{P}_{G}$. Let $\mathbb{F}$ be an algebraically closed field and consider the action of ${\rm Fun}(G,\mathbb{F}^{\times})$ on $\mathbb{A}_{S}$ given by
$\alpha\cdot(x_{\{g,h\}})_{\{g,h\}\in S}=\big{(}(d\alpha)(g,h)\,x_{\{g,h\}}\big{)}_{\{g,h\}\in S}.$
Then the vanishing ideal of the closure of the orbit of $\bar{1}_{S}$ in $\mathbb{A}_{S}$ is spanned by the binomials $r_{1}-r_{2}$ arising from all surviving higher-order identities, i.e., satisfying $r_{1}r_{2}^{-1}\in\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$. More generally, the vanishing ideal of the closure of the orbit of any $\varepsilon\in{\rm T}_{S}={\rm Fun}(S,\mathbb{F}^{\times})$ is spanned by the binomials $\varepsilon(r_{2})r_{1}-\varepsilon(r_{1})r_{2}$, where we make the identification ${\rm Fun}(S,\mathbb{F}^{\times})={\rm Hom}(\mathcal{F}_{S},\mathbb{F}^{\times})$.
Proof.
The orbit of $\bar{1}_{S}$ is the range of the function $f_{S}\colon{\rm Fun}(G,\mathbb{F}^{\times})\to\mathbb{A}_{S}$, $\alpha\mapsto\alpha\cdot\bar{1}_{S}$. Since $\alpha\cdot\bar{1}_{S}=(d\alpha)\bar{1}_{S}$, the range of $f_{S}$ is precisely $B^{2}_{S}(\mathbb{F}^{\times})$ defined in Section 3. For any monomial $r$ in variables indexed by $i\in S$, denote by $\lambda_{r}$ the restriction of the corresponding polynomial function $\mathbb{A}_{S}\to\mathbb{F}$ to the torus ${\rm Fun}(S,\mathbb{F}^{\times})$. First we claim that $\lambda_{r_{1}}|_{B^{2}_{S}(\mathbb{F}^{\times})}=\lambda_{r_{2}}|_{B^{2}_{S}(\mathbb{F}^{\times})}$ if and only if $r_{1}r_{2}^{-1}\in\mathcal{\langle}\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$.
In the course of the proof we use the following notation: if $B$ is a subgroup of an abelian group $A$ (written multiplicatively), then
$B^{\perp}:=\{f\in{\rm Hom}(A,\mathbb{F}^{\times})\mid f(B)=1\}\subset{\rm Hom}(A,\mathbb{F}^{\times}).$
Similarly, if $H$ is a subgroup of ${\rm Hom}(A,\mathbb{F}^{\times})$, then
$H^{\perp}:=\{a\in A\mid\chi(a)=1\mbox{ for all }\chi\in H\}\subset A.$
It is clear that $\lambda_{r_{1}}|_{B^{2}_{S}(\mathbb{F}^{\times})}=\lambda_{r_{2}}|_{B^{2}_{S}(\mathbb{F}^{\times})}$ if and only if the function $\lambda_{r_{1}}/\lambda_{r_{2}}$ maps $B^{2}_{S}(\mathbb{F}^{\times})$ to $1$, and the latter condition is equivalent to $r_{1}r_{2}^{-1}\in B^{2}_{S}(\mathbb{F}^{\times})^{\perp}\subset\mathcal{F}_{S}$.
Theorem 19 and Corollary 21 give us that the kernel of the restriction map ${\rm Hom}(\mathcal{Z}_{S},\mathbb{F}^{\times})\to{\rm Hom}(\mathcal{K}_{S},\mathbb{F}^{\times})$ is $B^{2}_{S}(\mathbb{F}^{\times})$. Since $\mathcal{Z}_{S}=\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ and $\mathcal{K}_{S}=(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})/\langle\mathcal{R}_{S}\rangle$, the kernel of the restriction map ${\rm Hom}(\mathcal{F}_{S},\mathbb{F}^{\times})\to{\rm Hom}(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S},\mathbb{F}^{\times})$ is also $B^{2}_{S}(\mathbb{F}^{\times})$. By definition, this kernel is $(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp}$, so we get $B^{2}_{S}(\mathbb{F}^{\times})=(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp}$ and hence $B^{2}_{S}(\mathbb{F}^{\times})^{\perp}=((\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp})^{\perp}$. Finally, $\mathcal{F}_{S}/(\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})\simeq\mathcal{I}_{S}$ is a free abelian group, so its points are separated by homomorphisms to $\mathbb{F}^{\times}$. Therefore, $((\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S})^{\perp})^{\perp}=\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$, which completes the proof of the claim.
Let $I$ be the vanishing ideal of the closure of $B^{2}_{S}(\mathbb{F}^{\times})$ in $\mathbb{A}_{S}$. Since $B^{2}_{S}(\mathbb{F}^{\times})$ is dense in its closure, we have
$I=\{\varphi\in\mathbb{F}[\mathbb{A}_{S}]\mid\varphi|_{B^{2}_{S}(\mathbb{F}^{\times})}=0\}.$
A binomial $r_{1}-r_{2}$ arises from a surviving higher-order identity if and only if $r_{1}r_{2}^{-1}\in\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$. It follows from the above claim that all these binomials $r_{1}-r_{2}$ belong to $I$. Modulo a linear combination of such binomials, any $\varphi\in\mathbb{F}[\mathbb{A}_{S}]$ is congruent to a linear combination of monomials $r_{j}$ such that $r_{i}r_{j}^{-1}\notin\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ for $i\neq j$. But the restrictions $\lambda_{r_{j}}|_{B^{2}_{S}(\mathbb{F}^{\times})}$ are pairwise distinct multiplicative characters of $B^{2}_{S}(\mathbb{F}^{\times})$, hence they are linearly independent over $\mathbb{F}$. Therefore the ideal $I$ is spanned by the binomials $r_{1}-r_{2}$ arising from surviving higher-order identities. This finishes the proof of the first claim of the theorem
The second part of theorem’s statement follows, since the ${\rm T}_{S}$-orbit of $\varepsilon$ is ${\rm B}^{2}_{S}(\mathbb{F}^{\times})\varepsilon$. ∎
Remark 28.
The group $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ of surviving higher-order identities is a free abelian group of rank $\leqslant|S|$, hence the vanishing ideal of the closure of the orbit of any $\varepsilon\in{\rm T}_{S}$ is generated by the binomials $\varepsilon(r_{2})r_{1}-\varepsilon(r_{1})r_{2}$, where $r_{1}r_{2}^{-1}$ ranges over a generating set of $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$. Moreover, when the group $G$ is finite, a generating set of $\langle\mathcal{R}_{G}\rangle\cap\mathcal{F}_{S}$ can be efficiently computed.
Corollary 29.
Let $G$ be a finite abelian group.
(1) The vanishing ideal of $X(\bar{1})$ is spanned by the binomials $r_{1}-r_{2}$ arising from all higher-order identities.
(2) For $S\in\mathcal{S}(G)$, we have $X(\bar{1}_{S})\subset X(\bar{1})$ if and only if, for any higher-order identity $r_{1}=r_{2}$, if $r_{1}$ involves only variables in $S$ then so does $r_{2}$ (no “weak violations” in the terminology of [39]).
(3) For any $\varepsilon\in X_{S}$ and $\varepsilon^{\prime}\in X$, we have $X(\varepsilon^{\prime})\subset X(\varepsilon)$ if and only if $S(\varepsilon^{\prime})\subset S$ and $\varepsilon^{\prime}\varepsilon^{-1}$ satisfies all surviving higher-order identities $r_{1}=r_{2}$ for $S$, i.e., arising from $r_{1}r_{2}^{-1}\in\langle R_{G}\rangle\cap\mathcal{F}_{S}$.
Theorem 27 is relevant when applying a $G$-graded contraction $\varepsilon$ to a finite-dimensional Lie algebra $\mathfrak{L}$ with a $G$-grading $\Gamma\colon\mathfrak{L}=\bigoplus_{g\in G}\mathfrak{L}_{g}$. Let $S:={\rm Supp}_{2}(\Gamma)$, i.e., the set
$\{\{g,h\}\in\mathcal{P}_{G}\mid[\mathfrak{L}_{g},\mathfrak{L}_{h}]\neq 0\}\subset\mathcal{P}_{{\rm Supp}(\Gamma)}.$
Then the Lie bracket $\mu$ of $\mathfrak{L}$ is the direct sum of $\mu_{g,h}\colon\mathfrak{L}_{g}\times\mathfrak{L}_{h}\to\mathfrak{L}_{gh}$, and the bracket of the contracted algebra $\mathfrak{L}^{\varepsilon}$ is obtained by multiplying each $\mu_{g,h}$ by $\varepsilon(g,h)$. Moreover, the action of $D(\Gamma)$ on $\mu$ from Section 2.2 corresponds to the action of ${\rm Fun}(G,\mathbb{F}^{\times})$ on $\bar{1}_{S}$. Thus the bracket of $\mathcal{L}^{\varepsilon}$ is in the closure of the $D(\Gamma)$-orbit of $\mu$ if and only if $\varepsilon$ is in the closure of the ${\rm Fun}(G,\mathbb{F}^{\times})$-orbit of $\bar{1}_{S}$. Theorem 27 implies the following:
Corollary 30.
The contracted algebra $\mathcal{L}^{\varepsilon}$ is a graded degeneration of $\mathcal{L}$ if and only if $\varepsilon$ satisfies all surviving higher-order identities for $S={\rm Supp}_{2}(\Gamma)$. In particular, if ${\rm Supp}_{2}(\Gamma)=\mathcal{P}_{G}$, then the graded degenerations are precisely the graded contractions in $X(\bar{1})$.
5 Graded contractions as lax monoidal functors
Now we will consider the way to formalize the “generic” point of view on contractions using the functors introduced by Moody and Patera in [28], which we will define in a more general setting of lax monoidal functors on the category ${\rm Vec}^{G}$ of $G$-graded vector spaces. As noted in Section 2.2, there are different types of equivalences for graded contractions, two of which correspond to isomorphisms in ${\rm Vec}^{G}$, namely, strong equivalence (Definition 10) and equivalence via normalization (Definition 11). One of them implies the other, and Weimar-Woods conjectured that they, in fact, coincide:
Conjecture 31 ([39]).
Two graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ are strongly equivalent, $\varepsilon\approx\varepsilon^{\prime}$, if and only if they are equivalent via normalization, $\varepsilon\sim_{\rm n}\varepsilon^{\prime}$.
Draper et al. recently showed that this conjecture holds for the grading group $\mathbb{Z}_{2}^{3}$, by considering a special $\mathbb{Z}_{2}^{3}$-grading on the simple Lie algebra of type $G_{2}$, see discussion in [9, Section 1]. Note that, if we assume that the isomorphism in Definition 10 holds for a specific $G$-graded Lie algebra $\mathfrak{L}$, rather than for all Lie algebras, then this statement is false, even if $\mathfrak{L}$ is semisimple and the support of the grading is generates $G$, as discussed in [9, Example 2.12]. Since there exists a $G$-graded Lie algebra with the second support $\mathcal{P}_{G}$ (see Proposition 7), it is clear that $\varepsilon\approx\varepsilon^{\prime}$ implies $S(\varepsilon)=S(\varepsilon^{\prime})$.
Proposition 32.
For the group $\mathbb{Z}$, if $\varepsilon\approx\varepsilon^{\prime}$, then there exists $\lambda\colon\mathbb{Z}\to\mathbb{F}^{\times}$ such that
$\varepsilon^{\prime}(m,n)=\varepsilon(m,n)(d\lambda)(m,n)\quad\mbox{for all }m\neq n.$
Proof.
Consider the Witt algebra $\mathcal{W}$ with the canonical basis given by $L_{n}$, $n\in\mathbb{Z}$, and the Lie bracket given by
$[L_{m},L_{n}]=(m-n)L_{m+n},\quad\forall m,n\in\mathbb{Z}.$
The algebra $\mathcal{W}$ has a natural $\mathbb{Z}$-grading,
$\mathcal{W}=\bigoplus_{n\in\mathbb{Z}}\mathcal{W}_{n},\quad\mathcal{W}_{k}=\langle L_{k}\rangle\quad\forall k\in\mathbb{Z}.$
Let $\varepsilon$ and $\varepsilon^{\prime}$ be strongly equivalent $\mathbb{Z}$-graded contractions of $\mathcal{W}$. This means that there exists an isomorphism of graded Lie algebras, $\varphi\colon\mathcal{W}^{\varepsilon}\to\mathcal{W}^{\varepsilon^{\prime}}$. Since $\dim\mathcal{W}_{n}=1$, for any $n\in\mathbb{Z}$ there exists $\lambda_{n}\in\mathbb{F}^{\times}$ such that $\varphi(L_{n})=\lambda_{n}L_{n}$. The equation $\varphi([L_{m},L_{n}]_{\varepsilon})=[\varphi(L_{m}),\varphi(L_{n})]_{\varepsilon^{\prime}}$ then gives the result. ∎
It is natural to consider Weimar–Woods conjecture within the categorical framework, where the notion of isomorphism is replaced with that of natural isomorphism. In fact, Moody and Patera initiated such consideration in [28, Section 4] even before Conjecture 31 appeared in the literature. They associated to graded contractions some specific endofunctors of the category of $G$-graded Lie algebras and further observed that the functors corresponding to contractions that are equivalent via normalization are naturally isomorphic.
We extend this analysis by providing a more general context using monoidal categories and lax monoidal functors, which applies not only to Lie algebras but also to Lie coloralgebras, in particular, to Lie superalgebras.
Lax monoidal functors. Here and in what follows we adopt the notation from [14]: a monoidal category is given by a quintuple $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$, where $\mathcal{C}$ is a category, $\otimes$ denotes the tensor product bifunctor, the natural isomorphism $\alpha$ determines the associativity constraint, $\mathsf{1}\in\mathcal{C}$ is the tensor unit, and $\iota\colon\mathsf{1}\otimes\mathsf{1}\to\mathsf{1}$ is an isomorphism, which satisfy certain axioms. When the context is clear, we may refer to this quintuple simply as $\mathcal{C}$. Furthermore, the isomorphism $\iota$ gives rise to the unit constraints $r$ and $l$ [14, Section 2.2], which may be used for an alternative difinition of a monoidal category [14, Definition 2.2.8].
A monoidal functor from $\mathcal{C}$ to $\mathcal{C}^{\prime}$ is a pair $(F,J)$, where $F\colon\mathcal{C}\to\mathcal{C}^{\prime}$ is a functor and $J_{X,Y}\colon F(X)\otimes^{\prime}F(Y)\to F(X\otimes Y)$ is a natural isomorphism, such that $F(\mathsf{1})$ is isomorphic to $\mathsf{1}^{\prime}$ and the diagram (12) below commutes. The notion of a lax monoidal functor between monoidal categories generalizes that of a monoidal functor by relaxing the condition of isomorphism with just morphism as follows:
Definition 33.
Let $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$ and $(\mathcal{C}^{\prime},\otimes^{\prime},\mathsf{1}^{\prime},\alpha^{\prime},\iota^{\prime})$ be monoidal categories. A lax monoidal functor between them is a triple $(F,J,\epsilon)$, where $F$ functor $F\colon\mathcal{C}\to\mathcal{C}^{\prime}$ together with coherence morphisms
  1. 1.
    a morphism $\epsilon\colon\mathsf{1}^{\prime}\to F(\mathsf{1})$,
  2. 2.
    a natural transformation $J(X,Y)\colon F(X)\otimes^{\prime}F(Y)\to F(X\otimes Y)$
satisfying the following conditions
  1. a.
    (associativity) For all objects $X,Y,Z\in\mathcal{C}$ the following diagram commutes:
[…]
  2. b.
    (unitality) For all $X\in\mathcal{C}$ the following diagrams commute:
[…]
    where $l$, $l^{\prime}$, $r$ and $r^{\prime}$ denote the left and right unit constraints of the two monoidal categories $\mathcal{C}$ and $\mathcal{C}^{\prime}$, respectively.
A natural transformation between two lax monoidal functors is just a natural transformation repsecting the monoidal structure.
Definition 34.
Let $(F,J,\epsilon)$ and $(G,I,\kappa)$ be lax monoidal functors from the monoidal category $(\mathcal{C},\otimes,\mathsf{1},\alpha,\iota)$ to the monoidal category $(\mathcal{C}^{\prime},\otimes^{\prime},\mathsf{1}^{\prime},\alpha^{\prime},\iota^{\prime})$. A monoidal natural transformation $\nu$ from $(F,J,\epsilon)$ to $(G,I,\kappa)$ is a natural transformation $\nu_{X}\colon F(X)\to G(X)$ between the underlying functors that is compatible with the tensor product and the unit in that the following diagrams commute for all objects $X,Y\in\mathcal{C}$:
[…]
$G$-graded vector spaces and monoidal structures on the identity endofunctor. Let $G$ be a group and $\mathbb{F}$ be any field. Consider the category ${\rm Vec}^{G}$ of $G$-graded vector spaces over $\mathbb{F}$. Morphisms in this category are linear maps which preserve the $G$-degree. The tensor product $V\otimes W$ on this category is defined by the formula
$(V\otimes W)_{g}:=\bigoplus_{x,y\in G\colon xy=g}V_{x}\otimes W_{y}.$
The unit object $\mathsf{1}$ is defined by $\mathsf{1}_{e}=\mathbb{F}$ and $\mathsf{1}_{g}=0$ for $g\neq e$. The associator $\alpha$ and the isomorphism $\iota$ are inherited from vector spaces. This endows the category ${\rm Vec}^{G}$ with a structure of an $\mathbb{F}$-linear monoidal category.
The simple objects in the category ${\rm Vec}^{G}$ are one-dimensional vector spaces $\delta_{g}$ indexed by the elements of the group $G$. The $G$-grading on $\delta_{g}$ is given by $(\delta_{g})_{x}=\mathbb{F}$ if $x=g$ and $(\delta_{g})_{x}=0$ otherwise. For these objects, we have $\delta_{g}\otimes\delta_{h}\simeq\delta_{gh}$.
The identity endofunctor $\mathrm{Id}\colon{\rm Vec}^{G}\to{\rm Vec}^{G}$ can be given the structure of a monoidal functor as follows. Given a function $\sigma\colon G\times G\to\mathbb{F}^{\times}$, we take as $\varepsilon$ the identity of $\mathsf{1}$ define a natural isomorphism $J^{\sigma}_{V,W}\colon V\otimes W\to V\otimes W$,
$J^{\sigma}_{V,W}(v\otimes w)=\sigma(g,h)v\otimes w\quad\mbox{for all}\ v\in V_{g},\ w\in W_{h}\quad\mbox{and}\quad g,h\in G.$
Then diagram (12) commutes if and only if
$\sigma(g,hk)\sigma(h,k)=\sigma(k,gh)\sigma(g,h)\quad\mbox{for all}\quad g,h,k\in G,$
which means that $\sigma$ is a two-cocycle of $G$ with coefficients in $\mathbb{F}^{\times}$ with trivial $G$-action. We denote the monoidal functor $(\mathrm{Id},J^{\sigma})$ by $F_{\sigma}$.
An algebra object (not necessarily associative) in ${\rm Vec}^{G}$ is a pair $(A,\mu)$, where $A$ is an object and $\mu\colon A\otimes A\to A$ is a morphism in ${\rm Vec}^{G}$. Of course, this is equivalent to saying that $A$ is a $G$-graded algebra. The monoidal functor $F_{\sigma}=(\mathrm{Id},J^{\sigma})$ allows us to twist algebra objects as follows: $F_{\sigma}\colon(A,\mu)\mapsto(A,\mu_{\sigma})$, where $\mu_{\sigma}$ is the composition of $J^{\sigma}_{A,A}\colon A\otimes A\to A\otimes A$ and $F_{\sigma}(\mu)=\mu\colon A\otimes A\to A$. More explicitly,
$\mu_{\sigma}(a\otimes b)=\sigma(g,h)\mu(a\otimes b)\quad\mbox{for all}\ a\in A_{g},\ b\in A_{h}\quad\mbox{and}\quad g,h\in G.$
Since we are interested in algebras that satisfy certain identities, we will need an additional structure, namely, a symmetric braiding on ${\rm Vec}^{G}$. For a braiding to exist, $G$ has to be abelian. Let $\beta\colon G\times G\to\mathbb{F}^{\times}$ be a bicharacter of $G$, meaning that
$\beta(gh,k)=\beta(g,k)\beta(h,k)\quad\mbox{and}\quad\beta(g,hk)=\beta(g,h)\beta(g,k)\ \ \mbox{for all}\ \ g,h,k\in G.$
Then $\beta$ defines a braiding on the monoidal category ${\rm Vec}^{G}$ as follows. The natural isomorphism $c_{V,W}\colon V\otimes W\to W\otimes V$ is given by
$c_{V,W}(v\otimes w)=\beta(g,h)w\otimes v\quad\mbox{for all}\ v\in V_{g},\ w\in W_{h}\quad\mbox{and}\quad g,h\in G,$
which is symmetric (i.e., satisfies $c_{W,V}\circ c_{V,W}={\rm id}_{V\otimes W}$) if and only if $\beta$ is (multiplicatively) skew-symmetric:
$\beta(g,h)=\beta(h,g)^{-1}\quad\mbox{for all}\ g,h\in G.$
We denote the resulting symmetric braided category by $({\rm Vec}^{G},\beta)$. In this category, one can define any variety of algebras given by a set of multilinear polynomial identities by replacing the usual action of the symmetric group of degree $n$ on $A^{\otimes n}$ by the action arising from the braiding. In particular, Lie algebras in $({\rm Vec}^{G},\beta)$ are known as Lie coloralgebras with commutation factor $\beta$.
The functor $F_{\sigma}$ is a braided monoidal functor from $({\rm Vec}^{G},\beta)$ to $({\rm Vec}^{G},\beta^{\prime})$, where
$\beta^{\prime}(g,h):=\frac{\sigma(g,h)}{\sigma(h,g)}\beta(g,h)\quad\mbox{for all}\ g,h\in G.$
This is the basis of the so-called Scheunert’s trick of transforming Lie coloralgebras to Lie superalgebras, see more details about this subject in Scheunert’s seminal paper [36] and in [1, 2, 32, 33]. Note that if $\sigma$ is symmetric, i.e., $\sigma(g,h)=\sigma(h,g)$, then $F_{\sigma}$ is a braided monoidal endofunctor on $({\rm Vec}^{G},\beta)$. Hence if $A$ is an algebra in $({\rm Vec}^{G},\beta)$ that belongs to a variety defined by a set of multilinear polynomial identities, then $F_{\sigma}(A)$ belongs to the same variety. In particular, Lie algebras are mapped to Lie algebras.
Graded contractions as Lax monoidal structures on the identity endofunctor. We equip the identity endofunctor $\mathrm{Id}$ with a lax monoidal structure. For the morphism $\epsilon\colon\mathbb{F}\to\mathbb{F}$ we simply take the identity ${\rm id}_{\mathbb{F}}$. Given a function $\varepsilon\colon G\times G\to\mathbb{F}$, we define a natural transformation $J^{\varepsilon}_{V,W}\colon V\otimes W\to V\otimes W$,
$J^{\varepsilon}_{V,W}(v\otimes w)=\varepsilon(g,h)v\otimes w\quad\mbox{for all}\ v\in V_{g},\ w\in W_{h}\quad\mbox{and}\quad g,h\in G.$
Then diagram (12) commutes if and only if
$\varepsilon(g,hk)\varepsilon(h,k)=\varepsilon(k,gh)\varepsilon(g,h)\quad\mbox{for all}\quad g,h,k\in G.$
As above, we want the monoidal functor $F_{\varepsilon}:=(\mathrm{Id},J^{\varepsilon},{\rm id}_{\mathbb{F}})$ to map Lie algebras to Lie algebras, so we will further require $\varepsilon$ to be symmetric. These two conditions make $\varepsilon$ a generic contraction in the sense of Proposition 5.
Remark 35.
If we drop the symmetry condition on $\varepsilon$, the functor $F_{\varepsilon}$ is a braided lax monoidal functor from $({\rm Vec}^{G},\beta)$ to $({\rm Vec}^{G},\beta^{\prime})$ if and only if
$\beta^{\prime}(g,h)\sigma(h,g)=\beta(g,h)\sigma(g,h)\quad\mbox{for all}\ g,h\in G.$
Hence, $F_{\varepsilon}$ maps Lie algebras in $({\rm Vec}^{G},\beta)$ to Lie algebras $({\rm Vec}^{G},\beta^{\prime})$.
The following result shows that Conjecture 31 holds when the strong equivalence of contractions is replaced with “natural” strong equivalence.
Proposition 36.
The lax monoidal functors $F_{\sigma}=(\mathrm{Id},J_{\sigma})$ and $F_{\sigma^{\prime}}=(\mathrm{Id},J_{\sigma^{\prime}})$ are naturally isomorphic if and only if $\varepsilon\sim_{\rm n}\varepsilon^{\prime}$, i.e., $\varepsilon^{\prime}=\varepsilon\,d\alpha$ for some $\alpha\colon G\to\mathbb{F}^{\times}$.
Proof.
Assume that there exists $\alpha\colon G\to\mathbb{F}^{\times}$ satisfying
$\sigma(g,h)=\alpha(gh)^{-1}\alpha(g)\alpha(h)\sigma^{\prime}(g,h)$ (13)
For each $V\in{\rm Vec}^{G}$ define an isomorphism $\nu_{V}\colon V\to V$ via $\nu_{V}(v):=\alpha(g)v$ for all $v\in V_{g}$ and $g\in G$. Since any morphism $f\colon V\to W$ in ${\rm Vec}^{G}$ preserves $G$-degrees, the diagram
${V}$${W}$${V}$${W}$$\scriptstyle{f}$$\scriptstyle{\nu_{V}}$$\scriptstyle{\nu_{W}}$$\scriptstyle{f}$
commutes, hence $\nu$ is a natural isomorphism from the identity endofunctor to itself. We need to verify that $\nu$ respects the lax structure. The commutativity of the first diagram from Definition 34, applied to $F_{\sigma}$ and $F_{\sigma^{\prime}}$,
[…]
is equivalent to equation (13). The second diagram from Definition 34 commutes since $\epsilon$ in the identity.
Conversely, suppose we have a natural isomorphism $\nu\colon F_{\sigma}\to F_{\sigma^{\prime}}$ of lax monoidal functors. Consider the simple objects $\delta_{g}$ in ${\rm Vec}^{G}$. Since $\nu_{\delta_{g}}\colon\delta_{g}\to\delta_{g}$ is an isomorphism, there exists $\alpha(g)\in\mathbb{F}^{\times}$ such that $\nu_{\delta_{g}}(v)=\alpha(g)v$ for all $v\in\delta_{g}$. In view of this, the commutativity of the diagram
[…]
is equivalent to equation (13). ∎
6 Conclusion
Graded contractions of Lie algebras had been studied intensively over the past three decades, mostly from computational perspective. All contractions had been computed and classified for some specific graded Lie algebras of low dimension, see, e.g., [6, 9, 8, 20, 21] and references therein. To the best of our knowledge, only Moody and Patera in [28] and Weimar–Woods in [39] studied generic graded contractions from theoretical point of view. We continued this direction but using different methods.
More specifically, Moody and Patera studied graded contractions from the perspectives of semigroup cohomology and endofunctors of the category of Lie algebras, see more details in Sections 3 and 5. Weimar-Woods, on the other hand, developed a theory based on the properties of what she called “surviving defining relations” and “surviving higher-order identities”, seemingly unrelated to cohomology. This led to a classification theorem for generic graded contractions [39, Theorem 7.1] over the fields of complex and real numbers.
The works of Moody, Patera and Weimar–Woods inspired us to consider generic graded contractions from the perspectives of group cohomology, affine algebraic geometry and monoidal categories. These connections provided a wider context to the classification problem of generic graded contractions, allowing us to obtain streamlined proofs and generalize the results of [28] and [39].
Now we summarize the results of this paper. We began Section 3 with establishing in Corollary 17 that the group $\mathcal{F}_{G}/\langle\mathcal{R}_{G}\rangle$ is a free abelian group of rank $\leqslant|G|$ using an analogue of the bar resolution. In Remark 18, we discussed the freeness of the group $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ claimed in [39, Theorem 6.5], whose proof appears to be incorrect. Nevertheless, the numerical experiments carried out for low-order grading groups lead us to pose the following question:
Is it true that, for any $S\subset\mathcal{P}_{G}$, the abelian group $\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle$ is free?
Then we introduced the abelian group $H^{2}_{S}(A)={\rm Hom}(\mathcal{F}_{S}/\langle\mathcal{R}_{S}\rangle,A)/B^{2}_{S}(A)$, which in the case $A=\mathbb{F}^{\times}$ and $S\in\mathcal{S}(G)$ classifies the $\mathbb{F}$-valued contractions with support $S$ up to equivalence via normalization. In Theorem 19, we presented exact sequences that allow one to find the group $H^{2}_{S}(A)$. The special cases when $\mathbb{F}$ is an algebraically closed or a real closed field were presented in Corollaries 21 and 22, respectively. These corollaries were related to Theorem 7.1 from [39] in Remark 23, thereby providing a cohomological interpretation of the notion of sign invariants from [39].
Although we established the classification theorem (Theorem 19) for generic graded contractions, there is still one important question to address:
For any finite group $G$, develop an efficient algorithm for finding the set of supports of generic graded contractions, $\mathcal{S}(G)$.
We touched upon this problem in Remark 13 and in Section 4.
In Section 4, we studied the affine algebraic variety of $G$-graded contractions and its cellular decomposition. Using this decomposition, in Theorem 27 we described the vanishing ideal of the closure of the ${\rm T}_{S}$-orbit of an arbitrary generic graded contraction with support $S$. This allowed us to relate graded contractions to graded degenerations, introduced in Section 2.1, and to Inönü–Wigner contractions.
We revisited Conjecture 31 from [39] in Section 5. Interpreting generic $G$-graded contractions as lax monoidal structures on the identity endofunctor of the monoidal category of $G$-graded vector spaces, we establish a functorial version of this conjecture.
Acknowledgements. SDK is grateful to Dr. Alex Bihlo for the support of his research thanks to funding from the Canada Research Chairs program, the InnovateNL Leverage R&D Program and the NSERC Discovery Grant. MVK is supported by NSERC Discovery Grant 2018-04883.